### v4

# Full exceptional collections on Fano threefolds and groups generated by spherical twists on K3 surfaces

Abstract.
For a Fano threefold admitting a full exceptional collection of vector bundles of length four we show that all full exceptional collections consist of shifted vector bundles. We prove this via a detailed study of the group generated by spherical twists on an anticanonical divisor. For example, we prove that this group is free and provide explicit generators.
Key words and phrases:
Fano threefold, exceptional collection, K3 surface, group of derived autoequivalences
2020 Mathematics Subject Classification:
14J45, 14J28, 14F08
The second author is a senior researcher at the Research Foundation Flanders (FWO). This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 885203).
1. Introduction
Bondal and Polishchuk conjectured [11, Conjecture 2.2] that the action of $B_{n}\rightthreetimes{\mathbb{Z}}^{n}$, where $B_{n}$ is the $n$-strand braid group, by mutations and shifts is transitive on full exceptional collections of length $n$ for any triangulated category. This conjecture has been recently disproved by Chang, Haiden and Schroll [15] who found a series of partially wrapped Fukaya categories producing counterexamples. [^1] Inspecting these counter examples suggests including derived autoequivalences in addition to mutations and shifts to achieve transitivity. Restricting to bounded derived categories ${\mathcal{D}}^{b}(X)$ of smooth projective varieties $X$, the conjecture has been verified in a small number of special cases, namely, by Rudakov [47], [46] for ${\mathbb{P}}^{2}$ and ${\mathbb{P}}^{1}\times{\mathbb{P}}^{1}$, by Kuleshov and Orlov [32] for all del Pezzo surfaces and by Akira, Okawa and Uehara [24] for the second Hirzebruch surface $\Sigma_{2}$, but in general it is still widely open.
Now consider threefolds with a full exceptional collection of four vector bundles. There are four types of such threefolds, namely, the Fano varieties $X={\mathbb{P}}^{3},Q_{3},V_{5}$} and $V_{22}$, where only $V_{22}$ has non-trivial moduli (see Lemma 3.5 and §7 for more details). Since $\operatorname{rk}K_{0}(X)=4=\dim X+1$, it follows from [9] (see Proposition 3.3) that the full $B_{n}\rightthreetimes{\mathbb{Z}}^{n}$-orbit of the given full exceptional collection consists of full exceptional collections of vector bundles, up to shifts. A converse to this result was obtained by Polishchuk [44]: he showed that the action of $B_{4}\rightthreetimes{\mathbb{Z}}^{4}$ on full exceptional collections of shifted vector bundles is transitive. However, even for ${\mathbb{P}}^{3}$ it was unknown if these constitute all full exceptional collections. The following result settles this:
Theorem 1.1 (Theorem 8.6, Corollary 8.7).
Let $X$ be a threefold with a full exceptional collection of four vector bundles. Then:
  1. (1)
    all full exceptional collections on $X$ consist of shifted vector bundles;
  2. (2)
    the action of $B_{4}\rightthreetimes{\mathbb{Z}}^{4}$ by mutations and shifts is transitive and free.
Note that in addition to the transitivity conjectured by Bondal and Polishchuk, we also prove that the action is free. By Polishchuk’s result, (1) implies the transitivity part of (2). However, we actually prove (2) first using a new approach detailed below and then we deduce (1) from it.
We remark that Theorem 1.1(1) does not rule out the existence of exceptional objects in ${\mathcal{D}}^{b}(X)$ which are not vector bundles, nor does it yield a classification of exceptional vector bundles on $X$, since even for ${\mathbb{P}}^{3}$ it is unknown whether every exceptional vector bundle can be included in a full exceptional collection.
Remark 1.2.
A general result by Positselski [45] states that if $\operatorname{rk}K_{0}(X)=\dim X+1$ then every full strong exceptional collection (see Definition 3.1) on $X$ consists of shifts of vector bundles. So to prove Theorem 1.1 it would be enough to show that every full exceptional collection is strong. We do not know how to do this directly however.
Our proof of Theorem 1.1(2) starts from the following elementary observation:
Observation 1.3 (Lemma 3.16).
The restriction of an exceptional $E$ object on $X$ to a generic anticanonical divisor $Y\subset X$ is a spherical object $E\mid_{Y}$. Let $\phi(E):=T_{E\mid_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ be the associated spherical twist. When extended to sequences of objects in ${\mathcal{D}}^{b}(X)$, $\phi$ intertwines the action of the braid group $B_{n}$ on the set of full exceptional collections in ${\mathcal{D}}^{b}(X)$ and the Hurwitz action of $B_{n}$ on $(\operatorname{Aut}{\mathcal{D}}^{b}(Y))^{n}$.
Recall that if $G$ is a group then the Hurwitz action of $B_{n}$ on $G^{n}$ is given by (see §3.2):
$\sigma_{i}(g_{1},\ldots,g_{i},g_{i+1},\ldots,g_{n})=(g_{1},\ldots,g_{i}g_{i+1}g_{i}^{-1},g_{i},\ldots,g_{n})$
The surface $Y$ defined in Lemma 1.3 is a K3 surface with Picard rank $1$ and we will obtain Theorem 1.1(2) as a corollary of studying $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ and its subgroup $\mathcal{G}$} generated by spherical twists for such surfaces. To do this we use the ground breaking work of Bridgeland [13] and Bayer-Bridgeland [4]. One of the main results of [4, 13] (also building on [23, 31, 43]) is the existence of a short exact sequence of groups (see Theorem 3.20)
(1.1) $1\to\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})\to 1,$
where $\operatorname{Aut}^{+}\operatorname{H}^{+}(Y,{\mathbb{Z}})$ is the group of Hodge isometries of $\operatorname{H}^{*}(Y,{\mathbb{Z}})$ which preserve the orientation of the positive $4$-planes and $\operatorname{Aut}^{0}{\mathcal{D}}(Y)$ is the product of the group generated by the double shift $[2]$ and the group freely generated by $T_{S}^{2}$ where $S$ runs through the spherical vector bundles.
It is well-known that K3 surfaces of Picard rank $1$ have a single discrete invariant, the minimal degree $d=2\delta$ of an ample divisor. In this introduction we will assume $d\geq 4$ for simplicity and refer to the text for the case $d=2$. First note that the algebraic cycles provide us with a subgroup of rank $3$ of $\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$
${\mathcal{N}}(Y):=\operatorname{H}^{0}(Y,{\mathbb{Z}})\oplus\operatorname{Pic}(Y)\oplus\operatorname{H}^{4}(Y,{\mathbb{Z}})\cong{\mathbb{Z}}^{3}$
which is equipped with the Mukai quadratic form $q(r,l,s)=\delta l^{2}-rs$. We define $\operatorname{Aut}^{+}{\mathcal{N}}(Y)$ as the isometries of ${\mathcal{N}}(Y)$ which preserve the orientation of the positive 2-planes. It turns out that there is a well-defined restriction map
(1.2) $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})\rightarrow\operatorname{Aut}^{+}{\mathcal{N}}(Y)$
which is injective (assuming $d\geq 4$, see Lemma B.5). The following precise descriptions of the groups $\operatorname{Aut}^{+}{\mathcal{N}}(Y)$ and $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})$ are well known and go back to Dolgachev [18] and Kawatani [30]. For the benefit of the reader, we give independent proofs of these results in the Appendix (see §A, §B).
Proposition 1.4 ([18, Theorem 7.1, Remark 7.2.2], [30, Theorem 3.3], see also Remark A.6, Proposition B.6).
Assume $d\geq 4$. We have the isomorphisms
(1.3) $\operatorname{Aut}^{+}\operatorname{H}(Y,{\mathbb{Z}})/\{\pm 1\}$ $\cong F(\delta)$
$\operatorname{Aut}^{+}{\mathcal{N}}(Y)/\{\pm 1\}$ $\cong\Gamma^{+}_{0}(\delta)$
where $F(\delta)$ is the Fricke group and $\Gamma^{+}_{0}(\delta)$ is the Atkin-Lehner group. The group $F(\delta)$ is the extension of the congruence subgroup
$\Gamma_{0}(\delta):=\left\{{\footnotesize\begin{pmatrix}a&b\\ c&d\end{pmatrix}}\in\operatorname{PSL}_{2}({\mathbb{Z}})\mid c\cong 0\operatorname{mod}\delta\right\}\subset\operatorname{PSL}_{2}({\mathbb{R}})$
by the Fricke involution $s_{\delta}=\left(\begin{smallmatrix}0&-1\\ \delta&0\end{smallmatrix}\right)$ and $\Gamma^{+}_{0}(\delta)$ is the extension of $\Gamma_{0}(\delta)$ obtained by adjoining the “Atkin-Lehner elements”, of which $s_{\delta}$ is a special case (see §3.4).
One has $F(\delta)/\Gamma_{0}(\delta)\cong{\mathbb{Z}}/2{\mathbb{Z}}$ and $\Gamma^{+}_{0}(\delta)/\Gamma_{0}(\delta)=({\mathbb{Z}}/2{\mathbb{Z}})^{c}$ where $c$ is the number of distinct prime factors in $\delta$. Hence (1.2) is in general not surjective. A first counter example is given by $\delta=6$, i.e. $d=12$.
The groups that appear in (1.3) are so-called “Fuchsian groups”, that is, they are discrete subgroups of the group of orientation preserving isometries of the complex upper half-plane $\mathbb{H}$} on which they act by Möbius transformations $z\mapsto(az+b)/(cz+d)$}. The fact that the hyperbolic plane $\mathbb{H}$} appears naturally in this context was already observed by Kawatani in [31]. In fact the existence of the exact sequence (1.1) ultimately boils down to the existence of an action, originally conjectured in [13] and proved in [4], of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ on a simply connected space $\operatorname{Stab}^{\dagger}(Y)$, a distinguished component of the moduli space of stability conditions for ${\mathcal{D}}^{b}(Y)$. Kawatani observed that $\operatorname{Stab}^{\dagger}(Y)$ is closely related to $\mathbb{H}$}. Leveraging this together with (1.3) we obtain the following, which was already observed by Fan and Lai in [19] (see (1.2) on p.4).
Observation 1.5 (Lemma 6.3).
Let $Y$ be a K3 surface of Picard rank $1$ and degree $d=2\delta\geq 4$. Let $\Delta^{+}(Y)=\{(r,l,s)\in{\mathcal{N}}(Y)\mid q(r,l,s)=-1,r>0\}$ and let ${\mathbb{H}}^{0}$ be the complement in $\mathbb{H}$} of the points $(l\delta+i\sqrt{\delta})/(r\delta)$ for $(r,l,s)\in\Delta^{+}(Y)$. Then
$\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]\cong\pi_{1}([{\mathbb{H}}^{0}/F(\delta)])$
where ${\mathbb{H}}/F(\delta)]$ is the orbifold quotient of ${\mathbb{H}}^{0}$ by $F(\delta)$.
${\mathbb{H}}^{0}/F(\delta)]$ is an orbifold Riemann surface, so in particular its fundamental group is well understood (see §5.2). This yields a good understanding of the group $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$. A similar approach can be applied to subgroups of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$ that contain $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$ such as (the image of) the group $\mathcal{G}$} generated by spherical twists, which was introduced above. It turns out that the image $G$ of $\mathcal{G}$} in $F(\delta)$ acts freely on ${\mathbb{H}}^{0}$ (see Lemma 5.8(1)), so that ${\mathbb{H}}^{0}/G$ is a non-compact Riemann surface. Hence ${\mathcal{G}}=\pi_{1}([{\mathbb{H}}^{0}/G])$ is actually a free group! Elaborating on this leads to our next main result.
Theorem 1.6 (Theorem 6.5).
Let $Y$ be a K3 surface of Picard rank $1$. Then the subgroup of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ generated by spherical twists is freely generated by a collection of spherical twists associated to spherical vector bundles.
The actual result in the text is more detailed and in fact gives a precise recipe to find the generators of $\mathcal{G}$} in terms of a suitable fundamental domain for the action of $G$ on $\mathbb{H}$}. We apply this recipe explicitly for some low-degree K3 surfaces in §7, in particular for those that appear as anticanonical sections of the Fano threefolds ${\mathbb{P}}^{3},Q_{3},V_{5}$ and $V_{22}$ which were introduced above. In that case we obtain:
Theorem 1.7 (Theorem 7.1).
Let $X={\mathbb{P}}^{3},Q_{3},V_{5}$} or $V_{22}$. Let $Y\subset X$ be a generic anticanonical divisor which is a K3 surface of Picard rank $1$ and degree $4,6,10$ and $22$ respectively. Then the group ${\mathcal{G}}\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ generated by all spherical twists is isomorphic to the free group $F_{4}$ on $4$ generators which are the spherical twists along the spherical vector bundles obtained by restricting a full exceptional collection of vector bundles on $X$ to $Y$.
We obtain Theorem 1.1(2) from Theorem 1.7 and the fact that the product of the spherical twists associated to a full exceptional collection only depends on the Serre functor on $X$.
Proposition 1.8 (Proposition 3.13).
Let $(E_{1},\cdots,E_{n})$ be a full exceptional collection in a Fano variety $X$ and let $i:Y\hookrightarrow X$ be a smooth anticanonical divisor. Let $T_{i}:=T_{E_{i}\mid_{Y}}$. One has $T_{1}\dots T_{n}(-)=-\otimes i^{*}\omega_{X}[2]$ in $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$
Combining this with Theorem [40, Theorem 2.6(1)] (see Theorem 3.18) we are able to deduce that any exceptional collection can be mutated into one furnishing a chosen set of generators for $\mathcal{G}$} (see Lemma 8.5) which essentially finishes the proof of Theorem 1.1(2).
To end this introduction we note that Theorem 1.7 may suggest the optimistic conjecture that $\mathcal{G}$} is always finitely generated. This is far from the case, however, and in fact in Corollary 5.11 we give precise conditions for it to be true. In §7 we discuss the nice example $d=8$. In that case it turns out that $\mathcal{G}$} is freely generated by $T_{{\mathcal{O}}(nH)}$, $n\in{\mathbb{Z}}$, where $H$ is a generator of $\operatorname{Pic}(Y)$.
2. Acknowledgement
We thank Tom Bridgeland for promptly answering our questions concerning [4] and Daniel Huybrechts for pointing out important missing references in the first version of the paper. The first author is greatly indebted to Alexey Bondal for introducing her to the crucial results of [4].
3. Generalities
We work over the complex numbers. Unless otherwise specified $X$ is a smooth projective variety. We denote by $\mathop{\text{{coh}}}(X)$ the category of coherent sheaves on $X$ and by ${\mathcal{D}}^{b}(X):={\mathcal{D}}^{b}(\mathop{\text{{coh}}}(X))$ the bounded derived category of $\mathop{\text{{coh}}}(X)$. A Fano variety is a smooth projective variety $X$ whose anticanonical divisor $K_{X}$ is ample. By a K3 surface $Y$ we mean a complex algebraic surface with trivial canonical bundle $\omega_{Y}\simeq{\mathcal{O}}_{Y}$ and $\operatorname{H}^{1}(Y,{\mathcal{O}}_{Y})=0$. We refer the reader to [22] for a comprehensive account on K3 surfaces.
Definition 3.1.
An object $E\in{\mathcal{D}}^{b}(X)$ is exceptional if $\operatorname{Hom}^{i}(E,E)=0$ for any $i\neq 0$ and $\operatorname{Hom}^{0}(E,E)\cong{\mathbb{C}}$. An ordered collection ${\mathcal{E}}=(E_{0},\dots,E_{m})$ of objects of ${\mathcal{D}}^{b}(X)$ is called an exceptional collection if each $E_{i}$ is exceptional and $\operatorname{Hom}^{\bullet}(E_{j},E_{k})=0$ for all $j>k$. An exceptional collection $\mathcal{E}$} is full if ${\mathcal{D}}^{b}(X)$ is the smallest triangulated subcategory of ${\mathcal{D}}^{b}(X)$ containing $E_{0},\dots,E_{m}$. We say that $\mathcal{E}$} is strong if $\operatorname{Hom}^{k}(E_{i},E_{j})=0$ for any $k\neq 0$.
Definition 3.2.
Let $(E,F)$ be an exceptional pair. We define the left (respectively right) mutation of $(E,F)$ as $(L_{E}F,E)$ (respectively $(F,R_{F}E)$), where the objects $L_{E}F$ and $R_{F}E$ are given by the distinguished triangles
$\operatorname{Hom}^{\bullet}(E,F)\otimes E\to F\to L_{E}F$
$R_{F}E\to E\to\operatorname{Hom}^{\bullet}(E,F)^{\ast}\otimes F$
Proposition 3.3 (Bondal, [9, Assertion 2.3]).
The braid group $B_{n}$ acts on the set of exceptional collections in ${\mathcal{D}}^{b}(X)$ of length $n$ by mutations. More precisely, the standard generator $\sigma_{i}\in B_{n}$, $1\leq i\leq n-1$ acts by
$\sigma_{i}(E_{0},\dots,E_{n-1})=(E_{0},\dots,E_{i-2},L_{E_{i-1}}E_{i},E_{i-1},E_{i+1},\dots,E_{n-1})$
Proposition 3.4 (Bondal, [9, Assertion 9.2]).
Let $X$ be a smooth projective variety with $\operatorname{rk}K_{0}(X)=\dim X+1$. Then mutations of full exceptional collections of shifted sheaves in ${\mathcal{D}}^{b}(X)$ also consist of shifted sheaves.
Lemma 3.5.
Let $X$ be a smooth projective threefold. The following conditions are equivalent.
  1. (1)
    ${\mathcal{D}}^{b}(X)$ admits a full exceptional collection of length $4$ consisting of vector bundles.
  2. (2)
    $X$ is Fano and ${\mathcal{D}}^{b}(X)$ admits a full exceptional collection of length $4$.
  3. (3)
    $X$ is Fano of Picard rank $1$ and $h^{1,2}(X)=0$.
  4. (4)
    In the classification of [26] (see also [7]) $X$ is one of the following types of Fano threefolds: ${\mathbb{P}}^{3}$, the 3-dimensional quadric $Q_{3}$, $V_{5}$ or $V_{22}$ (only $V_{22}$ has non-trivial moduli).
Proof.
[…]
3.1. Spherical twists
Definition 3.6.
An object $S\in{\mathcal{D}}^{b}(X)$ is spherical if $\omega_{X}\otimes S\cong S$ and $\operatorname{Hom}^{i}(S,S)={\mathbb{C}}$ if $i\in\{0,\dim X\}$} and $\operatorname{Hom}^{i}(S,S)=0$ otherwise. The set of spherical objects in ${\mathcal{D}}^{b}(X)$ up to isomorphism is denoted by $\operatorname{Sph}(X)$.
Definition 3.7 (Seidel-Thomas, [49]).
A spherical object $S\in{\mathcal{D}}^{b}(X)$ gives rise to an autoequivalence $T_{S}\in\operatorname{Aut}{\mathcal{D}}^{b}(X)$ (referred to as spherical twist) such that for every $A\in{\mathcal{D}}^{b}(X)$ there is a distinguished triangle
$\operatorname{Hom}^{\bullet}(S,A)\otimes S\xrightarrow{ev}A\to T_{S}(A).$
Alternatively, $T_{S}$ can be defined as the Fourier-Mukai transform with kernel $\operatorname{cone}(S\boxtimes S^{\vee}\to{\mathcal{O}}_{\Delta})\in{\mathcal{D}}^{b}(X\times X)$.
We also need a more general notion of a spherical twist with respect to a spherical functor.
Definition 3.8.
Let $F:{\mathcal{B}}\to{\mathcal{D}}$ be a functor between triangulated categories with fixed dg-enhancements in the sense of [10] and assume that $F$ has a right adjoint $R$ and a left adjoint $L$. The twist $T_{F}$ and the cotwist $C_{F}$ with respect to $F$ are defined via the distinguished triangles
$FR\xrightarrow{\varepsilon^{R}}{\operatorname{id}}_{{\mathcal{D}}}\to T_{F}$
$C_{F}\to{\operatorname{id}}_{{\mathcal{B}}}\xrightarrow{\eta^{R}}RF,$
where $\varepsilon^{R}$ and $\eta^{R}$ are the counit and the unit of the adjunction $(F,R)$. We say that the functor $F$ is spherical if $T_{F}$ and $C_{F}$ are autoequivalences.
Remark 3.9.
To make this definition sensible one needs to assume in addition that $F,R$ and $L$ have lifts to quasi-functors and consider cones that descend from the dg level. We refer the reader to [1] for a comprehensive treatment. In the geometric context one can circumvent discussing dg-enhancements by considering Fourier-Mukai kernels.
Lemma 3.10 ([21, Lemma 8.21]).
Let $S$ be a spherical object and $\Phi\in\operatorname{Aut}{\mathcal{D}}^{b}(X)$ an autoequivalence. Then
$T_{\Phi(S)}\cong\Phi T_{S}\Phi^{-1}$
In particular, for any two spherical objects $S_{1},S_{2}$, one has $T_{T_{S_{1}}(S_{2})}\cong T_{S_{1}}T_{S_{2}}T_{S_{1}}^{-1}$.
Passing from $S$ to $T_{S}$ loses some information, but not much. For simplicity we state the following lemma for $\dim X\geq 2$.
Lemma 3.11.
Assume $\dim X\geq 2$ and let $S,S^{\prime}$ be spherical objects. Then $T_{S}\cong T_{S^{\prime}}$ if and only if $S\cong S^{\prime}[n]$ for some $n\in{\mathbb{Z}}$.
Proof.
We discuss the non-obvious direction. So assume $T_{S}\cong T_{S^{\prime}}$. We then have $T_{S^{\prime}}(S)=T_{S}(S)=S[1-\dim X]$. On the other hand, $T_{S^{\prime}}(S)$ fits into the distinguished triangle
$\operatorname{Hom}^{\bullet}(S,S^{\prime})\otimes S^{\prime}\xrightarrow{}S\xrightarrow{0}S[1-\dim X],$
where the right arrow is 0 because $\operatorname{Hom}(S,S[1-\dim X])=0$. This implies that $S$ is a summand of $\operatorname{Hom}^{\bullet}(S,S^{\prime})\otimes S^{\prime}$. Since both $S$ and $S^{\prime}$ are indecomposable, this yields that $S=S^{\prime}[n]$ for some $n\in{\mathbb{Z}}$. ∎
Now let $X$ be a smooth Fano variety and $i:Y\hookrightarrow X$ a smooth anticanonical divisor. If $E\in{\mathcal{D}}^{b}(X)$ is an exceptional object, then the object $i^{*}E\in{\mathcal{D}}^{b}(Y)$ is spherical (see [49, Proposition 3.15]).
Lemma 3.12.
If $(E,F)$ is an exceptional pair in ${\mathcal{D}}^{b}(X)$, then
(3.1) $\operatorname{Hom}_{X}^{\bullet}(E,F)\cong\operatorname{Hom}_{Y}^{\bullet}(i^{*}E,i^{*}F)$
Proof.
Observe that by the adjunction formula
$\operatorname{Hom}_{Y}(i^{*}E,i^{*}F[k])\cong\operatorname{Hom}_{X}(E,F\otimes i_{*}{\mathcal{O}}_{Y}[k])$
for any $k\in{\mathbb{Z}}$. Applying $\operatorname{Hom}_{X}(E,-)$ to the distinguished triangle
$F\otimes\omega_{X}\to F\to F\otimes i_{*}{\mathcal{O}}_{Y}\rightarrow$
(obtained from identifying $\omega_{X}\cong{\mathcal{O}}_{X}(-Y)$) we get the long exact sequence
$\dots\to\operatorname{Hom}_{X}(E,F[k])\to\operatorname{Hom}_{X}(E,F\otimes i_{*}{\mathcal{O}}_{Y}[k])\to\operatorname{Hom}_{X}(E,F\otimes\omega_{X}[k+1])\to\dots$
By Serre duality $\operatorname{Hom}_{X}(E,F\otimes\omega_{X}[k+1])\cong\operatorname{Hom}_{X}(F,E[\dim X-k+1])^{*}=0$ for any $k\in{\mathbb{Z}}$, since $(E,F)$ is an exceptional pair. Hence (3.1) follows. ∎
Now suppose $X$ has a full exceptional collection
${\mathcal{D}}^{b}(X)=\langle E_{1},\dots,E_{n}\rangle$}.
We denote $T_{k}:=T_{i^{*}E_{k}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. The following fact is well known to experts, but we include the proof for the reader’s convenience (see e.g. [14, Theorem 1.1] for a special case).
Proposition 3.13.
There is an isomorphism of endofunctors of ${\mathcal{D}}^{b}(Y)$
(3.2) $T_{1}\dots T_{n}(-)=-\otimes i^{*}\omega_{X}[2].$
Proof.
Let ${\mathcal{E}}:=\bigoplus_{k=1}^{n}E_{k}$ and $A:=\operatorname{REnd}_{X}({\mathcal{E}})$. Then the functor
(3.3) $\operatorname{RHom}_{X}({\mathcal{E}},-):{\mathcal{D}}^{b}(X)\to{\mathcal{D}}^{b}(A)$
is an equivalence of categories with quasi-inverse given by $\otimes_{A}^{\mathbb{L}}{\mathcal{E}}$ (see [9]). Denote $\overline{{\mathcal{E}}}:=\bigoplus_{k=1}^{n}i^{*}E_{k}$. By [2, Theorem 3.0.1] and Lemma 3.12 the composition $T_{1}\dots T_{n}$ is the twist along the spherical functor
$\otimes_{A}^{\mathbb{L}}\overline{{\mathcal{E}}}:{\mathcal{D}}^{b}(A)\to{\mathcal{D}}^{b}(Y).$
Taking the composition with the equivalence (3.3) one observes that $T_{1}\dots T_{n}$ is also the twist along the spherical functor $i^{*}:{\mathcal{D}}^{b}(X)\to{\mathcal{D}}^{b}(Y)$. Thus, for any ${\mathcal{F}}\in{\mathcal{D}}^{b}(Y)$ we have
$T_{1}\dots T_{n}({\mathcal{F}})=T_{i^{*}}({\mathcal{F}})=\operatorname{cone}(i^{*}i_{*}{\mathcal{F}}\to{\mathcal{F}})={\mathcal{F}}\otimes{\mathcal{O}}_{Y}(-Y)[2],$
where the last equality is given by e.g. [21, Corollary 11.4]. ∎
Remark 3.14.
We will also consider collections of spherical objects which occur in the following way. Let $Y$ be a K3 surface of degree $2$. Then $Y$ is realised as a double covering $\pi:Y\to{\mathbb{P}}^{2}$ branched along a sextic curve. If $E\in{\mathcal{D}}^{b}({\mathbb{P}}^{2})$ is exceptional, then ${\mathbb{R}}\pi^{*}(E)\in{\mathcal{D}}^{b}(Y)$ is spherical ([49, Proposition 3.13, Example 3.14b]). One also checks that (3.2) holds with the right-hand side replaced by $\otimes i^{*}\omega_{X}[1]$.
3.2. The Hurwitz action
Definition 3.15.
Let $G$ be a group. Then the $n$-strand braid group $B_{n}$ with the standard generators $\sigma_{1},\dots,\sigma_{n-1}$ acts on $G^{n}$ via
$\sigma_{i}(g_{1},\ldots,g_{i},g_{i+1},\ldots,g_{n})$ $=(g_{1},\ldots,g_{i}g_{i+1}g_{i}^{-1},g_{i},\ldots,g_{n}).$
This action is referred to as the Hurwitz action. [^2] In [40] a different convention for the Hurwitz action is used, but this is inconsequential.
For $X$ Fano, $i:Y\hookrightarrow X$ a smooth anticanonical divisor and $E$ an exceptional object in ${\mathcal{D}}^{b}(X)$ we denote by $\phi(E):=T_{i^{*}(E)}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ the corresponding spherical twist. We extend $\phi$ to full exceptional collections in $X$ by applying it term wise.
Lemma 3.16.
$\phi$ intertwines the action of the braid group $B_{n}$ on the set of full exceptional collections in ${\mathcal{D}}^{b}(X)$ defined in Proposition 3.3 and the Hurwitz action of $B_{n}$ on $(\operatorname{Aut}{\mathcal{D}}^{b}(Y))^{n}$.
Proof.
It suffices to prove that for an exceptional pair $(E,F)$ in ${\mathcal{D}}^{b}(X)$ we have
$\phi(L_{E}F)=\phi(E)\phi(F)\phi(E)^{-1}.$
By Lemma 3.10 one has
$\phi(E)\phi(F)\phi(E)^{-1}=T_{T_{i^{*}(E)}(i^{*}(F))}.$
Hence we need to show that $i^{*}(L_{E}F)=T_{i^{*}(E)}(i^{*}(F))$ or, equivalently,
$i^{*}\operatorname{cone}(\operatorname{Hom}^{\bullet}_{X}(E,F)\otimes E\to F)\cong\operatorname{cone}(\operatorname{Hom}^{\bullet}_{Y}(i^{*}E,i^{*}F)\otimes i^{*}E\to i^{*}F)$
This now follows from Lemma 3.12. ∎
Let $F_{m}=\langle x_{1},\dots,x_{m}\rangle$ be the free group on $m$ generators. We are going to use the following results concerning the Hurwitz action on powers of free groups.
Theorem 3.17 (e.g. [17, XI, Corollary 1.8]).
The stabiliser of $(x_{1},\dots,x_{m})$ under the Hurwitz action of $B_{m}$ on $(F_{m})^{m}$ is trivial.
Theorem 3.18 ([40, Theorem 2.6(1)]).
Under the Hurwitz action of $B_{n}$ on $(F_{m})^{n}$, the orbit of $t=(t_{1},\dots,t_{n})$ with $t_{i}\in\{x_{1},\dots,x_{m}\}$ is given by the set of elements in
$\bigcup_{\tau\in S_{n}}C(t_{\tau(1)})\times\cdots\times C(t_{\tau(n)})$
whose product is $t_{1}\cdots t_{n}$, where $C(x)$ denotes the conjugacy class of $x$.
3.3. Stability conditions and derived autoequivalences of K3 surfaces
Let $Y$ be a K3 surface with Picard rank $1$. Let $H$ be an ample generator of $\operatorname{Pic}(Y)\cong\operatorname{NS}(Y)$} and put $H\cdot H=d:=2\delta$. The extended Neron-Severi group (or equivalently the numerical Grothendieck group)
${\mathcal{N}}(Y)=\operatorname{H}^{0}(Y,{\mathbb{Z}})\oplus\operatorname{Pic}(Y)\oplus\operatorname{H}^{4}(Y,{\mathbb{Z}})\cong{\mathbb{Z}}^{3}$
is equipped with the Mukai pairing given by
$\langle(r_{1},l_{1},s_{1}),(r_{2},l_{2},s_{2})\rangle:=dl_{1}l_{2}-r_{1}s_{2}-r_{2}s_{1}$
or, equivalently, a quadratic form $q(r,l,s)=\delta l^{2}-rs$. In particular ${\mathcal{N}}(Y)$ has signature $(2,1)$. For an object $E\in{\mathcal{D}}^{b}(Y)$ let
(3.4) $v(E):=(\operatorname{rk}(E),c_{1}(E),c_{1}(E)^{2}/2-c_{2}(E)+\operatorname{rk}(E))\in{\mathcal{N}}(Y)$
be the corresponding Mukai vector. Then for $E,F\in{\mathcal{D}}^{b}(Y)$ we have
$\langle v(E),v(F)\rangle=-\chi(E,F)=-\sum_{i\in{\mathbb{Z}}}(-1)^{i}\dim\operatorname{Hom}^{i}(E,F).$
Hence if $E\in{\mathcal{D}}^{b}(Y)$ is a $2$-spherical object and $v(E)=(r,l,s)$, then
$\langle v(E),v(E)\rangle=dl^{2}-2rs=-2$
We say that a Mukai vector $\rho\in{\mathcal{N}}(Y)$ is spherical (or a root) if $\langle\rho,\rho\rangle=-2$. Set
$\Delta(Y):=\{\rho\in{\mathcal{N}}(Y)\ |\ \langle\rho,\rho\rangle=-2\}$
$\Delta^{+}(Y):=\{\rho=(r,l,s)\in\Delta(Y)\ |\ r>0\}$
For any $\rho\in\Delta^{+}(Y)$ there exists a unique spherical vector bundle $E_{\rho}$ with $v(E_{\rho})=\rho$ and any spherical sheaf on $Y$ is a slope stable vector bundle. This is a combination of the results of Yoshioka [55] and Mukai [39], see footnote (2) on page 10 of [4].
If $E\in{\mathcal{D}}^{b}(Y)$ is a spherical object, then the spherical twist $T_{E}$ acts on cohomology by the reflection at the hyperplane orthogonal to $v(E)\in{\mathcal{N}}(Y)$:
$s_{v(E)}(\alpha)=\alpha+\langle\alpha,v(E)\rangle v(E).$
Let $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})\subset\operatorname{Aut}\operatorname{H}^{*}(Y,{\mathbb{Z}})$ be the subgroup of all Hodge isometries which preserve the orientation of the positive 4-planes. By a result due to Huybrechts, Macrì and Stellari ([23, Theorem 2], see also [22, Corollary 16.3.13]) the subgroup $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})$ is precisely the image of the natural map $\operatorname{Aut}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}\operatorname{H}^{*}(Y,{\mathbb{Z}})$}. We denote the kernel of this map by $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$. The space
(3.5) ${\mathcal{P}}(Y):=\{v\in{\mathcal{N}}(Y)_{{\mathbb{C}}}\ |\ \operatorname{Re}(v),\operatorname{Im}(v)\text{ span a positive 2-plane}\}$
has two connected components. Let ${\mathcal{P}}^{+}(Y)$ be the component containing $\exp(iH)=(1,i,-\delta)$. Define an open subset
${\mathcal{P}}_{0}^{+}(Y):={\mathcal{P}}^{+}(Y)\setminus\bigcup_{\rho\in\Delta}\rho^{\perp},$
where $\rho^{\perp}$ denotes the orthogonal complement with respect to the extended Mukai pairing on ${\mathcal{N}}(Y)_{\mathbb{C}}$.
For the details regarding stability conditions on K3 surfaces and in general we refer the reader to [12], [13]. Recall that a numerical stability condition $(Z,{\mathcal{P}})$ on ${\mathcal{D}}^{b}(Y)$ consists of a group homomorphism $Z:{\mathcal{N}}(Y)\to{\mathbb{C}}$ (referred to as the central charge) and a collection of full (abelian) subcategories ${\mathcal{P}}(\phi)\subset{\mathcal{D}}^{b}(Y)$ satisfying a set of axioms, for all phases $\phi\in{\mathbb{R}}$. A stability condition is called locally finite if each abelian category ${\mathcal{P}}(\phi)$ has finite length. Denote the space of numerical locally finite stability conditions on ${\mathcal{D}}^{b}(Y)$ by $\operatorname{Stab}(Y)$. As shown in [12, Corollary 1.3], $\operatorname{Stab}(Y)$ is a finite-dimensional complex manifold. Recall that there is a natural left action of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ on $\operatorname{Stab}(X)$. This action is faithful and commutes with the right action of the universal cover $\widetilde{\operatorname{GL}^{+}(2,{\mathbb{R}})}$ of the group of orientation-preserving automorphisms $\operatorname{GL}^{+}(2,{\mathbb{R}})$ ([12, Lemma 8.2]). There is a subgroup ${\mathbb{C}}\subset\widetilde{\operatorname{GL}^{+}(2,{\mathbb{R}})}$ acting freely by rotating the central charge and relabelling the phases.
Since the Mukai pairing is non-degenerate, for every $Z:{\mathcal{N}}(Y)\to{\mathbb{C}}$ there exists $Z^{\vee}\in{\mathcal{N}}(Y)_{{\mathbb{C}}}$ such that $Z(-)=\langle Z^{\vee},$-$\rangle$. Let $\pi:\operatorname{Stab}(Y)\to{\mathcal{N}}(Y)_{{\mathbb{C}}}$ be the forgetful map sending a stability condition $(Z,{\mathcal{P}})$ to $Z^{\vee}$ (by [12, Theorem 1.2], this is a local homeomorphism).
Theorem 3.19 (Bridgeland, [13, Theorem 1.1]).
Let $Y$ be a K3 surface. There exists a connected component $\operatorname{Stab}^{\dagger}(Y)$ of $\operatorname{Stab}(Y)$ which is mapped onto ${\mathcal{P}}_{0}^{+}(Y)$ by $\pi$. Moreover, the induced map $\pi:\operatorname{Stab}^{\dagger}(Y)\to{\mathcal{P}}_{0}^{+}(Y)$ is a Galois covering and the subgroup of $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$ preserving $\operatorname{Stab}^{\dagger}(Y)$ can be identified with the group of deck transformations of $\pi$.
Theorem 3.19 implies that there is a map $\varphi:\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\to\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$. It was conjectured by Bridgeland ([13, Conjecture 1.2]) that $\operatorname{Stab}^{\dagger}(Y)$ is simply-connected and is preserved by $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ (hence, in particular, $\varphi$ is bijective). For K3 surface with Picard rank $1$ this conjecture was proved by Bayer and Bridgeland:
Theorem 3.20 (Bayer-Bridgeland, [4, Theorem 1.3]).
Let $Y$ be a K3 surface with Picard rank $1$. Then $\operatorname{Stab}^{\dagger}(Y)$ is contractible. In particular, there is an exact sequence of groups
(3.6) $1\to\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)\cong\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\to\operatorname{Aut}{\mathcal{D}}^{b}(Y)\to\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})\to 1.$
Moreover, for any $\rho\in\Delta(Y)$, the group $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$ acts transitively on the set of all spherical objects with Mukai vector $\rho$.
The following explicit description of the fundamental group $\pi_{1}({\mathcal{P}}_{0}^{+}(Y))$ was obtained by Kawatani:
Theorem 3.21 (Kawatani, [31, Theorem 1.3]).
Let $Y$ be a K3 surface with Picard rank 1. The group $\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\cong{\mathbb{Z}}\times\langle T_{E_{\rho}}^{2}\ |\ \rho\in\Delta^{+}(Y)\rangle$ is the product of $\mathbb{Z}$} acting by even shifts and the free group generated by the squares $T_{E_{\rho}}^{2}$ of spherical twists corresponding to all spherical vector bundles $E_{\rho}$, $\rho\in\Delta^{+}(Y)$.
Remark 3.22.
We will use these explicit generators of $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$ later on. Note that in general the fundamental group of a topological space does not have a canonical set of generators, and the fundamental groups associated with different base points are not canonically identified, so let us explain this briefly. One can show that ${\mathcal{P}}_{0}^{+}(Y)$ is isomorphic to $\operatorname{GL}_{2}^{+}({\mathbb{R}})\times{\mathbb{H}}^{0}$ (see §4.1 below), where ${\mathbb{H}}^{0}$ is the upper half plane $\mathbb{H}$} minus a discrete set of points corresponding to roots in $\Delta^{+}(Y)$. Hence $\pi_{1}({\mathcal{P}}_{0}^{+}(Y))\cong{\mathbb{Z}}\times\pi_{1}({\mathbb{H}}^{0})$. Let ${\mathcal{V}}(Y)\subset{\mathbb{H}}^{0}$ be the open subset obtained by removing the vertical line segments connecting the real line and the punctures (see Figure 1 on page 11 of [4]). ${\mathcal{V}}(Y)$ is open and contractible and has a canonical lift to $\operatorname{Stab}^{\dagger}(Y)$, as explained in [13, §6]. Its points are stability conditions corresponding to tiltings of the heart $\mathop{\text{{coh}}}(Y)$. Pick an arbitrary base point $y$ in ${\mathcal{V}}(Y)$ (since ${\mathcal{V}}(Y)$ is contractible the choice does not matter) and consider a collection of paths connecting it to punctures and not intersecting the relative interiors of the vertical segments. The corresponding collection of loops around the punctures in the direction of the orientation (they intersect the vertical segments exactly once) gives a canonical set of generators of $\pi_{1}({\mathbb{H}}^{0},y)$.
Corollary 3.23.
Let $Y$ be a K3 surface with Picard rank 1. Let ${\mathcal{G}}\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ the subgroup generated by spherical twists and let $G$ be its image in $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$. The map
$\operatorname{Sph}(Y)/({\mathbb{Z}}[2]\times{\mathcal{G}})\rightarrow\Delta(Y)/G$
$S\mapsto v(S)$
is a bijection.
Proof.
Surjectivity follows from the fact that every $\rho\in\Delta^{+}(Y)$ is the Mukai vector of the spherical vector bundle $E_{\rho}$. To prove injectivity assume $S,T\in\operatorname{Sph}(Y)$ are such that $f(v(S))=v(T)$ for $f\in G$. Lifting $f$ to $F\in{\mathcal{G}}$ we obtain $v(F(S))=v(T)$. But then by Theorem 3.21 we have $T=F^{\prime}(F(S))$ for $F^{\prime}\in{\mathbb{Z}}[2]\times\{H^{2}\mid H\in{\mathcal{G}}\}\subset{\mathbb{Z}}[2]\times{\mathcal{G}}$}. This finishes the proof. ∎
3.4. Fuchsian groups
We refer to [29] for the basics on Fuchsian groups. The group $\operatorname{PSL}_{2}({\mathbb{R}})$ acts on the upper half plane model $\mathbb{H}$} of the hyperbolic plane by Möbius transformations, i.e.
$z\mapsto\frac{az+b}{cz+d}$
with
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{PSL}_{2}({\mathbb{R}})$
[…]
Definition 3.24.
A Fuchsian group is a discrete subgroup of $\operatorname{PSL}_{2}({\mathbb{R}})$.
Theorem 3.25 ([29, Theorem 2.2.6, Corollary 2.2.7]).
Let $\Gamma\subset\operatorname{PSL}_{2}({\mathbb{R}})$ be a subgroup. Then the following are equivalent:
  1. (1)
    $\Gamma$ is Fuchsian.
  2. (2)
    $\Gamma$ acts properly discontinuously on $\mathbb{H}$} (that is, for any compact set $K\subset{\mathbb{H}}$ there are only finitely many $g\in\Gamma$ with $gK\cap K\neq\varnothing$).
  3. (3)
    All orbits $\Gamma x$ for $x\in{\mathbb{H}}$ are discrete subsets of $\mathbb{H}$}.
Theorem 3.25 implies that for a Fuchsian group $\Gamma$ a non-trivial $\Gamma$-stabiliser of a point $x\in{\mathbb{H}}$ is a maximal finite cyclic subgroup (elliptic subgroup). Conversely, one can show that every elliptic subgroup of $\Gamma$ is the stabiliser of a point $x\in{\mathbb{H}}$. We say that an element $g\in\Gamma$ is elliptic if $\langle g\rangle$ is an elliptic subgroup. A parabolic element $g\in\Gamma$ is a non-trivial element stabilising a point $x\in{\mathbb{R}}\cup\{\infty\}$. Theorem 3.25 implies in particular that the topological quotient $V:={\mathbb{H}}/\Gamma$ is well defined and is an orbifold Riemann surface. The orbifold points of $V$ correspond to non-congruent elliptic elements of $\Gamma$ and the cusps correspond to non-congruent parabolic elements. The orbifold fundamental group $\pi_{1}(V)$ is isomorphic to $\Gamma$. The following lemma says that the quotient is not pathological.
Lemma 3.26.
Let $\Gamma$ be a Fuchsian group. The orbifold points are discrete in ${\mathbb{H}}/\Gamma$.
Proof.
This follows from the fact that the set of orbifold points is $\Gamma$-invariant and discrete in $\mathbb{H}$} (see [29, Corollary 2.2.8]). ∎
An important piece of data associated with a Fuchsian group is a fundamental domain (or region).
Definition 3.27 ([29, §3.1]).
Let $\Gamma$ be a Fuchsian group. A closed region $\mathcal{F}$$\subset{\mathbb{H}}$ (i.e. the closure of a non-empty open set $\mathcal{F}$$^{\circ}$, the interior of $\mathcal{F}$}) is said to be a fundamental domain for $\Gamma$ if
  1. (i)
    $\bigcup_{g\in\Gamma}{\mathcal{F}}={\mathbb{H}}$.
  2. (ii)
    ${\mathcal{F}}^{\circ}\cap g({\mathcal{F}}^{\circ})=\varnothing$ for all $g\in\Gamma-\{1\}$.
A convex fundamental domain always exists, see [29, §3.2]. In the sequel we will need the following result. We define a translation $t_{\alpha}$ for $\alpha\in{\mathbb{R}}$ to be a Möbius transformation of the form $z\mapsto z+\alpha$.
Lemma 3.28.
Let $\Gamma$ be a Fuchsian group which is normalised by some translation $t_{\alpha}$, $\alpha>0$ such that $\langle\Gamma,t_{\alpha}\rangle$ is still Fuchsian. Then $\Gamma$ has a convex fundamental domain containing $\infty$ in its closure in $\overline{{\mathbb{H}}}$.
Proof.
For $g=\left(\begin{smallmatrix}a&b\\ c&d\end{smallmatrix}\right)$, $\det g=1$ define
$E_{g}=\{z\in{\mathbb{H}}\mid|cz+d|>1\}$
Note that $E_{g}$ is a half plane in the hyperbolic plane, and in particular it is convex. Assume first that some $t_{\alpha}$ is actually contained in $\Gamma$ and let $\alpha$ be minimal with this property. Let $\Gamma_{0}$ be the subgroup of $\Gamma$ generated by $t_{\alpha}$. Then according to [50, Definition 5.6] we can take $\mathcal{F}$} to be the Ford domain
${\mathcal{F}}=\overline{\Sigma_{c,\alpha}\cap\bigcap_{g\in\Gamma-\Gamma_{0}}E_{g}}$
with
$\Sigma_{c,\alpha}=\{z\in{\mathbb{H}}\mid c<\operatorname{Re}z<c+\alpha\}$
for $c\in{\mathbb{R}}$ arbitrary.
Assume now that $\Gamma$ does not contain a non-trivial translation and put $\bar{\Gamma}=\langle\Gamma,t_{\alpha}\rangle$. Then $\bar{\Gamma}=\coprod_{n\in{\mathbb{Z}}}t_{\alpha}^{n}\Gamma$ and in particular $\bar{\Gamma}-\bar{\Gamma}_{0}=\{t_{\alpha}^{n}g\mid n\in{\mathbb{Z}},g\in\Gamma-\{1\}\}$. Note that $E_{t_{\alpha}^{n}g}=E_{g}$. It follows that $\bar{\Gamma}$ has a fundamental domain
$\overline{\Sigma_{c,\alpha}\cap\bigcap_{g\in\Gamma-\{1\}}E_{g}}$
From the coset decomposition and Lemma 3.29 below we then get that $\Gamma$ has a fundamental domain which may be called a Ford domain of infinite width:
$\bigcup_{n}t^{n}_{\alpha}(\overline{\Sigma_{c,t}\cap\bigcap_{g\in\Gamma-\{1\}}E_{g}})=\overline{\bigcap_{g\in\Gamma-\{1\}}E_{g}}$
Here we have used that the union is obviously locally finite and furthermore that $t_{\alpha}(E_{g})=E_{t_{\alpha}gt^{-1}_{\alpha}}$ so that $\bigcap_{g\in\Gamma-\{1\}}E_{g}$ is $t_{\alpha}$-invariant. ∎
We have used the following standard lemma.
Lemma 3.29.
Let $\Lambda\subset\Gamma$ be Fuchsian groups and let $\Gamma=\coprod_{i\in I}\Lambda\gamma_{i}$ be a coset decomposition. Assume that $\mathcal{F}$} is a fundamental domain for $\Gamma$ such that the collection of closed sets $(\gamma_{i}{\mathcal{F}})_{i}$ is locally finite. Then $\bigcup_{i\in I}\gamma_{i}{\mathcal{F}}$ is a fundamental domain for $\Lambda$.
Proof.
See e.g. [53, Theorem 12]. The proof in loc. cit. assumes that $I$ is finite, but one checks that is goes through under the local finiteness hypothesis. ∎
Recall the standard definition of the Hecke congruence subgroups, which are examples of Fuchsian groups: [^3] We take our congruence subgroups to be subgroups of $\operatorname{PSL}_{2}({\mathbb{R}})$ rather than $\operatorname{SL}_{2}({\mathbb{R}})$ since we are mainly interested in their action on the upper half plane $\mathbb{H}$} by Möbius transformations.
$\Gamma_{0}(n):=\bigg\{\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\operatorname{PSL}_{2}({\mathbb{Z}})\ |\ c\equiv 0\pmod{n}\bigg\}$
Definition 3.30.
[…]
$\begin{pmatrix}ae&b\\ cn&de\end{pmatrix}\in M_{2}({\mathbb{Z}}).$
It is a trivial exercise that an Atkin-Lehner associated to $e$ exists if and only if $e$ is an exact divisor of $n$, i.e. $\gcd(e,n/e)=1$. The set of images of all Atkin-Lehner elements is the Atkin-Lehner group and is denoted by $\Gamma^{+}_{0}(n)\subset\operatorname{PGL}_{2}^{+}({\mathbb{Q}})$.
The group $\Gamma_{0}^{+}(n)$ contains $\Gamma_{0}(n)$, and for square free $n$, it is its normaliser. We have $\Gamma^{+}_{0}(n)/\Gamma_{0}(n)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{c}$ where $c$ the number of distinct prime factors in $n$. The groups $\Gamma_{0}^{+}(n)$ and the corresponding modular curves $X_{0}^{+}(n):={\mathbb{H}}/\Gamma_{0}^{+}(n)$ have been studied extensively. For instance, Ogg [41] observed that the primes $p$ for which $X_{0}^{+}(p)$ has genus 0 are precisely the prime factors of the order of the Monster group.
The special Atkin-Lehner element
$s_{n}=\begin{pmatrix}0&-1\\ n&0\end{pmatrix}\in\Gamma_{0}^{+}(n)$
is called the Fricke involution.
It is clear that $s_{n}$ normalises $\Gamma_{0}(n)$. The Fricke group is defined as
$F(n):=\langle\Gamma_{0}(n),s_{n}\rangle\subseteq\Gamma_{0}^{+}(n)$
and one has $F(n)/\Gamma_{0}(n)={\mathbb{Z}}/2{\mathbb{Z}}$. So $F(n)=\Gamma_{0}^{+}(n)$ if and only if $n$ is a prime power. Hence the smallest $n$ for which $F(n)$ and $\Gamma^{+}_{0}(n)$ differ is $n=6$. For example, we have
$\begin{pmatrix}2&-1\\ 6&-2\end{pmatrix},\begin{pmatrix}3&-1\\ 12&-3\end{pmatrix}\in\Gamma_{0}^{+}(6)\setminus F(6).$
4. Geometric considerations around ${\mathcal{P}}^{+}(Y)$, ${\mathcal{P}}_{0}^{+}(Y)$
4.1. Introduction
In this section we give a more or less self-contained discussion of some results of Kawatani in [31]. Note that ${\mathcal{P}}(Y)$, as introduced in §3.3 is equipped with obvious commuting $O(q,{\mathbb{R}})$ and $\operatorname{GL}_{2}^{+}({\mathbb{R}})$-actions (the latter is free and is obtained via identifying ${\mathbb{C}}\cong{\mathbb{R}}^{2}$). We let $O^{+}(q,{\mathbb{R}})$ be the index two subgroup of $O(q,{\mathbb{R}})$ which preserves the connected component ${\mathcal{P}}^{+}(Y)$.
Following [13] Kawatani considers the morphism
$e:{\mathbb{H}}\rightarrow{\mathcal{P}}^{+}(Y):z\mapsto\exp(zH)$
and shows that this defines a $\operatorname{GL}_{2}^{+}({\mathbb{R}})$-equivariant isomorphism
${\mathbb{H}}\times\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong{\mathcal{P}}^{+}(Y)$
[…]
${\mathcal{P}}_{0}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong{\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})^{0}_{<0}\cong{\mathbb{H}}^{0}.$
4.2. The hyperboloid model
Let $W$ be a real vector space of dimension n+1{n+1} equipped with a quadratic form $q$ of signature $(+\cdots+,-)$. Let $\langle-,$-$\rangle$ be the corresponding pairing. In particular, $\langle x,x\rangle=2q(x)$. Put
$\operatorname{Gr}(2,W)_{>0}:=\{V\subset W\mid\dim V=n,q\ {\mid}\ V\text{ is positive definite}\}.$
By sending $V$ to $V^{\perp_{q}}$ and using the fact that $q$ has signature $(n,1)$ we see that [^5] We use the convention that ${\mathbb{P}}(W)=(W-\{0\})/{\mathbb{R}}^{\ast}$.
$\operatorname{Gr}(2,W)_{>0}\cong{\mathbb{P}}(W)_{<0}:=\{\bar{x}\in{\mathbb{P}}(W)\mid q(x)<0\}.$
Consider
$W_{-1}:=\{x\in W\mid q(x)=-1\}.$
If $p_{0}\in W_{-1}$ then the restriction $\langle p_{0},-\rangle\mid W_{-1}\neq 0$ (since $\langle p_{0},x\rangle=0$ implies $x\in p_{0}^{\perp}$ and hence $q(x)>0$, because of the signature of $q$). Thus the sign of the function $\langle p_{0},$-$\rangle$ divides $W_{-1}$ into two connected components which are exchanged by $x\mapsto-x$ (it is easy to see that there are precisely two components). Let $W_{-1}^{+}$ be the component which contains $p_{0}$ (i.e. $\langle p_{0},-\rangle\mid W_{-1}^{+}<0$). Then
(4.1) $W_{-1}^{+}\rightarrow{\mathbb{P}}(W)_{<0}:x\mapsto\bar{x}$
is a diffeomorphism. Moreover one can check that the Lorentzian metric $ds^{2}=q(dx)$ on $W$ restricts to a Riemannian metric on $W_{-1}$. The pair $(W_{-1}^{+},q(dx)\mid W_{-1}^{+})$ is the so-called hyperboloid model for hyperbolic space. The hyperboloid metric on $W_{-1}$ may be transferred to ${\mathbb{P}}(W)_{<0}$ via (4.1). It is obvious from the construction that $O(q,{\mathbb{R}})$ acts by isometries on $W_{-1}$ and ${\mathbb{P}}(W)_{<0}$.
4.3. Application to K3 surfaces
Now let $Y$ be a K3 surface with Picard rank $1$. We are going to use the notations and definitions from §3.3. Recall from §3.3 that the quadratic form induced by the Mukai pairing on the extended Néron-Severi group ${\mathcal{N}}(Y)$ is $q(r,l,s)=\delta l^{2}-rs$, where $d=2\delta$ is the degree of $Y$. For $z\in{\mathbb{C}}$ we put $e(z):=\exp(zH)\in{\mathcal{N}}(X)_{\mathbb{C}}$. If $z=x+iy$ then we have
(4.2) $e(x+iy)$ $=1+(x+iy)H+\frac{1}{2}(x+iy)^{2}H^{2}$
$\cong(1,xH,\frac{x^{2}-y^{2}}{2}H^{2})+i(0,yH,xyH^{2})$
$\cong(1,x,\delta(x^{2}-y^{2}))+i(0,y,2\delta xy).$
Note that if $z\in{\mathbb{C}}$ then it follows from (4.2) that
(4.3) $\begin{gathered}\langle\operatorname{Re}e(z),\operatorname{Im}e(z)\rangle=0,\\ q(\operatorname{Re}e(z))=q(\operatorname{Im}e(z))=\delta y^{2}.\end{gathered}$
It follows from (4.3) that $e(z)\in{\mathcal{P}}(Y)$ for $z\in{\mathbb{H}}$, where $\mathbb{H}$} denotes the upper half-plane and ${\mathcal{P}}(Y)$ is defined in (3.5). If we identify ${\mathbb{C}}\cong{\mathbb{R}}^{2}$ then it is clear that
[…]
Let ${\mathcal{N}}(Y)_{{\mathbb{R}},-1}^{+}$ be the connected component of ${\mathcal{N}}(Y)_{{\mathbb{R}},-1}$ containing the points such that $r>0$.
Lemma 4.1.
The composition
(4.4) ${\mathbb{H}}\xrightarrow{e}{\mathcal{P}}^{+}(Y)\rightarrow{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong\operatorname{Gr}(2,{\mathcal{N}}(Y)_{\mathbb{R}})_{>0}\cong{\mathcal{N}}(Y)_{{\mathbb{R}},-1}^{+}$
is given by
(4.5) ${\mathbb{H}}\rightarrow{\mathcal{N}}(Y)_{{\mathbb{R}},-1}^{+}:x+iy\mapsto\frac{1}{\sqrt{\delta}y}(1,x,\delta(x^{2}+y^{2})).$
Proof.
As explained in §4.2 the composition (4.4) sends $z$ to $w$ where $w\in{\mathbb{R}}^{3}$ is orthogonal to $\operatorname{Re}e(z),\operatorname{Im}e(z)$ and which satisfies q(w)=1q(w)=-1, $r>0$. Using (4.2) one checks that one may take $w$ as in (4.5). ∎
Remark 4.2.
One can check that if we equip ${\mathcal{N}}(Y)^{+}_{{\mathbb{R}},-1}$ with the Riemannian metric introduced in §4.2 then the pullback of this metric under (4.5) is the standard metric on $\mathbb{H}$} given by
$ds^{2}=\frac{dx^{2}+dy^{2}}{y^{2}}.$
Corollary 4.3.
The composition
${\mathbb{H}}\xrightarrow{e}{\mathcal{P}}^{+}(Y)\rightarrow{\mathcal{P}}^{+}(Y)/\operatorname{GL}^{+}_{2}({\mathbb{R}})\cong\operatorname{Gr}(2,{\mathcal{N}}(Y)_{\mathbb{R}})_{>0}\cong{\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$
is given by
(4.6) ${\mathbb{H}}\rightarrow{\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}:(x,y)\mapsto[1:x:\delta(x^{2}+y^{2})].$
The inverse map is given by
(4.7) $p:{\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}\rightarrow{\mathbb{H}}:[r:l:s]\mapsto\frac{l+i\sqrt{\delta^{-1}rs-l^{2}}}{r}.$
Proof.
To verify the second formula we compute for
$x=\frac{l}{r},\quad y=\frac{\sqrt{\delta^{-1}rs-l^{2}}}{r}$
$\delta(x^{2}+y^{2})$ $=\frac{\delta l^{2}}{r^{2}}+\frac{rs-\delta l^{2}}{r^{2}}=\frac{s}{r}$
so that
$1:x:\delta(x^{2}+y^{2})]=[r:l:s]\qed$
Corollary 4.4.
If $\rho:=[r:l:s]\in{\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ corresponds to a root, i.e. if $\delta l^{2}-rs=-1$, then
(4.8) $p(\rho)=\frac{l\delta+i\sqrt{\delta}}{r\delta}$
4.4. Transfer of the $\operatorname{PSL}_{2}({\mathbb{R}})$ action
Let $q$ be as in §4.3. In this section we discuss how to transfer the action of $\operatorname{PSL}_{2}({\mathbb{R}})$ on $\mathbb{H}$} to ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$.
Lemma 4.5.
Let $\phi:{\mathbb{H}}\rightarrow{\mathbb{H}}$ be the Möbius transformation corresponding to a matrix
$\begin{pmatrix}a&b\\ c&d\end{pmatrix}$.
Then we have a commutative diagram
${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$$\mathbb{H}$${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$$\mathbb{H}$$\scriptstyle{\Phi}$$\scriptstyle{p}$$\scriptstyle{\phi}$$\scriptstyle{p}$
where $\Phi$ is given by
(4.9) $\begin{pmatrix}r\\ l\\ s\end{pmatrix}\mapsto\begin{pmatrix}d^{2}&2cd&c^{2}/\delta\\ bd&bc+ad&ac/\delta\\ b^{2}\delta&2ab\delta&a^{2}\end{pmatrix}\begin{pmatrix}r\\ l\\ s\end{pmatrix}$
Proof.
We will prove that $\Phi\circ p^{-1}=p^{-1}\circ\phi$. If $z=x+iy$ then the formula (4.6) can be written as
$p^{-1}(z)=[1:(z+\bar{z})/2:\delta z\bar{z}]$
We then compute
$(\Phi\circ p^{-1})(z)$ $=\begin{pmatrix}d^{2}&2cd&c^{2}/\delta\\ bd&bc+ad&ac/\delta\\ b^{2}\delta&2ab\delta&a^{2}\end{pmatrix}\begin{pmatrix}1\\ (\bar{z}+z)/2\\ \delta z\bar{z}\end{pmatrix}$
$=\begin{pmatrix}(cz+d)(c\bar{z}+d)\\ ((az+b)(c\bar{z}+d)+(a\bar{z}+b)(cz+d))/2\\ \delta(az+b)(a\bar{z}+b)\end{pmatrix}$
$=(p^{-1}\circ\phi)(z)\qed$
The subgroup of index two in $O(q,{\mathbb{R}})$ that preserves ${\mathcal{N}}(Y)_{{\mathbb{R}},-1}^{+}$ is denoted by $O^{+}(q,{\mathbb{R}})$. It is the group of isometries of the hyperboloid model. The subgroup $\operatorname{SO}^{+}(q,{\mathbb{R}})$ is the group of orientation preserving isometries of the hyperboloid model.
Note that in the natural map
(4.10) $\operatorname{Aut}^{+}({\mathcal{N}}(Y)_{\mathbb{R}})\rightarrow O^{+}(q,{\mathbb{R}})\cong\operatorname{Isom}({\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0})$
the two ++’s have different meanings. In fact (4.10) kills 1-1 (the image of $[1]$) and its image is $\operatorname{SO}^{+}(q,{\mathbb{R}})$. More precisely (4.10) induces an isomorphism
(4.11) $\operatorname{Aut}^{+}({\mathcal{N}}(Y)_{\mathbb{R}})/\{\pm 1\}\rightarrow\operatorname{SO}^{+}(q,{\mathbb{R}})$
The subgroup of orientation preserving isometries of $\mathbb{H}$} is $\operatorname{PSL}_{2}({\mathbb{R}})$. One checks that $\det\Phi=(\det\phi)^{3}$. So from Lemma 4.1 and Lemma 4.5 we obtain isomorphisms of Lie groups
(4.12) $\operatorname{PSL}_{2}({\mathbb{R}})\cong\operatorname{SO}^{+}(q,{\mathbb{R}}):\phi\mapsto\frac{\Phi}{\det\phi}$
We state a few results without proof. They can be verified easily using Lemma 4.1.
Corollary 4.6.
When $\alpha\in{\mathbb{Z}}$ the isometry on ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ induced by $\otimes{\mathcal{O}}_{Y}(\alpha)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ is given by the formula.
(4.13) $t_{\alpha}:{\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}\rightarrow{\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}:[r:l:s]\mapsto[r:l+\alpha r:s+2\delta\alpha l+\delta\alpha^{2}r]$
(note that this formula makes sense for $\alpha\in{\mathbb{R}}$). $t_{\alpha}$ corresponds to
$t_{\alpha}:{\mathbb{H}}\rightarrow{\mathbb{H}}:z\mapsto z+\alpha.$
Corollary 4.7.
Let $S\in{\mathcal{D}}^{b}(Y)$ be a spherical object with $v(S)=(r,l,s)\in\Delta^{+}(Y)$ a positive root (cfr (3.4)). Recall that the action of $T_{S}$ on ${\mathbb{P}}({\mathcal{N}}(Y)_{{\mathbb{R}}})_{<0}$ is given by
$s_{v(S)}:v\mapsto v+v(S)\langle v(S),v\rangle.$
Then $s_{v(S)}$ corresponds to the point reflection at $p([r:l:s])=\frac{l}{r}+\frac{i}{r\sqrt{\delta}}\in{\mathbb{H}}$. Explicitly it is given by the Möbius transformation
$\begin{pmatrix}\delta l&-s\\ \delta r&-\delta l\end{pmatrix}\in F(\delta).$
5. Groups generated by spherical reflections in $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y)$
5.1. Free products generated by elements of order two
Lemma 5.1.
Let $G$ be a group and assume that $R\subset G-\{1\}$ are order two elements such that
$G=\ast_{r\in R}\langle r\rangle,$
where $\ast$ denotes the free product. Then any element of order two in $G-\{1\}$ is conjugate to some element in $R$.
Proof.
Let $x\in G-\{1\}$ be such that $x^{2}=1$. Expressing $x$ as a reduced word in $R$ we see that $x$ must be palindromic. If $x$ has odd length then it is of the form $wrw^{-1}$. Even length is impossible since then $x=ww^{-1}=1$. ∎
Lemma 5.2.
Let $G$ be a group and assume that $R\subset G-\{1\}$ are order two elements such that
$G=\ast_{r\in R}\langle r\rangle$
Let $S\subset G-\{1\}$ be a collection of elements such that
  1. (a)
    $G$ is generated by $S$;
  2. (b)
    For each $s\in S$ we have $s^{2}=1$;
  3. (c)
    $S$ is closed under conjugation.
Then
  1. (1)
    $R$ is a set of orbit representatives for the action of $G$ on $S$ by conjugation (in particular $R\subset S$).
  2. (2)
    $S$ is the set of order two elements in $G-\{1\}$.
Proof.
Let $s\in S$. By Lemma 5.1 $s$ is conjugate to some $r\in R$ and by (c) $r\in S$. So every element in $S$ is conjugate to an element in $R\cap S$. Assume first that $R\cap S\neq R$ and choose $r\in R\setminus R\cap S$. Since by (a) $S$ generates $G$ it follows that $r$ is a product of conjugates of elements in $R\cap S$. Hence in $G/[G,G]$, $\bar{r}$ is a linear combination of $\bar{r}^{\prime}$ for $r^{\prime}\in S\cap R$. However this is impossible since $\bar{r}^{\prime}$ for $r^{\prime}\in R$ is a ${\mathbb{Z}}/2{\mathbb{Z}}$-basis for $G/[G,G]$. We conclude that $R\subset S$ and furthermore every conjugation orbit in $S$ intersects $R$. If the intersection has cardinality $\geq 2$ then we find two elements in $R$ than are conjugate. Passing again to $G/[G,G]$ shows that this is impossible. So $R$ is a set of orbit representatives. This finishes the proof of (1).
If $g\in G-\{1\}$ is any element of order two then by Lemma 5.1 it is conjugate to an element in $R$. Hence since $R\subset S$, it is in $S$. This finishes the proof of (2). ∎
Lemma 5.3.
Let the notations and assumptions be as in Lemma 5.2. We have a presentation
(5.1) $1\rightarrow F\rightarrow{\mathcal{G}}\rightarrow G\rightarrow 0$
where $\mathcal{G}$} is defined as the free group generated by symbols $\tilde{r}$ for $r\in R$ and $F$ is the free group generated by $w\tilde{r}^{2}w^{-1}$ where $w$ is a square free product of the $\tilde{r}$, $r\in R$. In particular, sending $w\tilde{r}^{2}w^{-1}\mapsto\bar{w}r\bar{w}^{-1}\in S$ provides a natural indexation of the given free generators of $F$ by the elements of $S$.
Proof.
The fact that the given elements are free generators of $F$ follows naturally from the topological proof of the Nielsen-Schreier theorem (see [52, §2.2.3]). The coset graph for ${\mathcal{G}}/F$ is a tree with doubled edges and contracting the natural spanning tree provides our generators. We leave the details to the reader. ∎
Remark 5.4.
The given generators of $F$ are not canonical in any way. For example there is an uncountable number of spanning trees one may contract, each yielding a different set of generators.
5.2. Preliminaries on orbifold Riemann surfaces
Let $V$ be an orbifold Riemann surface with a discrete set of orbifold points $(x_{i})_{i}$ with orders $(e_{i})_{i}$. Let $V^{\circ}=V-\{(x_{i})_{i}\}$. So $V^{\circ}$ is an ordinary Riemann surface with punctures given by $(x_{i})_{i}$. Choose a non-orbifold base point $x\in V$. For any path in $\gamma:x\rightarrow x^{\prime}$ we denote by $\tau_{\gamma}$ the small loop around $x^{\prime}$, starting in $x$, in the direction of the orientation, obtained by thickening $\gamma$.
Lemma 5.5.
Choose non-intersecting paths $\gamma_{i}$ from $x$ to $x_{i}$ in $V^{\circ}$ and let $({{\tau}}_{i})_{i}=(\tau_{\gamma_{i}})_{i}$ be the corresponding loops around the punctures. Then we have a presentation
(5.2) $\pi_{1}(V,x)=\pi_{1}(V^{\circ},x)/(({{\tau}}_{i}^{e_{i}})_{i}).$
Proof.
See the discussion on the construction of $\pi_{1}(V,x)$ in [48, p423ff.]. ∎
Corollary 5.6.
Let $V$, $(x_{i})_{i}$, $(e_{i})_{i}$, ${{\tau}}_{i}$, $x$ be as above. Let $V$ be an orbifold Riemann surface with a Let $W$ be the coarse moduli space of $V$ (i.e. $W$ is obtained from $V$ by replacing the orbifold points by ordinary points). If $W$ is contractible then there is a presentation
$\pi_{1}(V,x)=\ast_{i}(\langle{{\tau}}_{i}\rangle/({{\tau}}_{i}^{e_{i}}))$
Proof.
Since $W$ is contractible, by the uniformization theorem we have $W={\mathbb{H}}$ or $W={\mathbb{C}}$. We have $V^{\circ}=W-\{(x_{i})_{i}\}$ and it easy to see that $\pi_{1}(V^{\circ},x)=\ast_{i}\langle{{\tau}}_{i}\rangle$. We now use (5.2). ∎
Remark 5.7.
Assume $V=[{\mathbb{H}}/G]$ is an orbifold quotient for a Fuchsian group $G$. Let $(y_{i})_{i}\in{\mathbb{H}}$ be orbit representatives for the fixed points of elements of $G$. Assume the $(y_{i})_{i}$ are contained in a connected fundamental domain $\mathcal{F}$} and choose $y$ in the interior of $\mathcal{F}$} (the $(y_{i})_{i}$ must be on the boundary). Let $x$, $(x_{i})_{i}$ be the images of $y$, $(y_{i})_{i}$ in $V$. Then a system of paths $(\gamma_{i})_{i}$ from $x$ to $x_{i}$ as in Lemma 5.5 can be constructed as the image of a corresponding system of paths $\tilde{\gamma}_{i}$ from $y$ to $y_{i}$ in $\mathcal{F}$}. The $y_{i}$ are then the fixed points of the ${{\tau}}_{i}$ under the identification $G\cong\pi_{1}(V,x)$ corresponding to the choice of $y$.
When $V$ is of finite type, i.e. $\{(x_{i})_{i}\}=\{x_{1},\ldots,x_{m}\}$ and $V=\overline{V}-\{x_{m+1},\ldots,x_{m+n}\}$ with $\overline{V}$ compact of genus $g$, then (5.2) can be made more precise in the sense that one has the well-known presentation for $\pi_{1}(V,x)$ given by
$a_{1}b_{1}a_{1}^{-1}b_{1}^{-1}\cdots a_{g}b_{g}a_{g}^{-1}b_{g}^{-1}{{\tau}}_{1}\cdots{{\tau}}_{m}{{\tau}}_{m+1}\cdots{{\tau}}_{m+n}=1,\quad{{\tau}}_{i}^{e_{i}}=1,i=1,\ldots,m$
where we have augmented $\{{{\tau}}_{1},\ldots{{\tau}}_{n}\}$ with additional loops $\{\tau_{n+1},\ldots,\tau_{m+n}\}$, going around the punctures $x_{m+1},\cdots x_{m+n}$ (see [48, p423ff.]). The loops ${{\tau}}_{1},\ldots,{{\tau}}_{m+n}$ can be constructed as in Remark 5.7 by including points at infinity.
It follows that if $n\geq 1$ then $\pi_{1}(V,x)$ is a free product of cyclic groups
(5.3) $\langle a_{1}\rangle\ast\langle b_{1}\rangle\ast\cdots\ast\langle a_{g}\rangle\ast\langle b_{g}\rangle\ast\langle{{\tau}}_{m+1}\rangle\ast\cdots\ast\langle{{\tau}}_{m+n-1}\rangle\ast\langle{{\tau}}_{1}\rangle/({{\tau}}_{1}^{e_{1}})\ast\cdots\ast\langle{{\tau}}_{m}\rangle/({{\tau}}_{m}^{e_{m}})\\ \cong{\mathbb{Z}}^{\ast 2g+n-1}\ast{\mathbb{Z}}/e_{1}{\mathbb{Z}}\ast\cdots\ast{\mathbb{Z}}/e_{m}{\mathbb{Z}}.$
5.3. Groups generated by point reflections in the hyperbolic plane
Recall from Definition 3.30 that $\Gamma_{0}^{+}(\delta)\subset\operatorname{PSL}_{2}({\mathbb{R}})$, $\delta\in{\mathbb{N}}$ is the group consisting of all Atkin-Lehner elements.
Lemma 5.8.
Let $G$ be a subgroup of $\Gamma^{+}_{0}(\delta)$ which is generated by point reflections $(s_{i})_{i}$. Let $S$ be the union of the conjugacy classes of the $(s_{i})_{i}$. Then
  1. (1)
    The set of fixed points of elements of $G$ is the union of the orbits of the fixed points of the points $(s_{i})_{i}$. Moreover the stabilisers of fixed points are isomorphic to ${\mathbb{Z}}/2{\mathbb{Z}}$.
  2. (2)
    The Riemann surface ${\mathbb{H}}/G$ is contractible.
  3. (3)
    Every element of order two in $G$ is conjugate to some $s_{i}$.
  4. (4)
    There exists a set of orbit representatives $R\subset S$ such that $G$ is the free product of the ${\mathbb{Z}}/2{\mathbb{Z}}$ groups generated by the corresponding reflections.
Proof.
Let $V=[{\mathbb{H}}/G]$ be the orbifold so that the ordinary quotient $W={\mathbb{H}}/G$ is the topological quotient. Choose a base point $x$ and paths $({{\tau}}_{i})_{i}$ as in Lemma 5.5. Since $\mathbb{H}$} is simply connected, after choosing a lift $y\in{\mathbb{H}}$ of $x$, we have a canonical identification $G\cong\pi_{1}(V,x)$. Since every path in $W$ can be lifted to a path in $V$, after a small deformation, we have a surjection:
$G\cong\pi_{1}(V,x)\rightarrow\pi_{1}(W,x).$
Since $W$ is an ordinary Riemann surface, its fundamental group has no torsion ([20, Proposition IV.6.5]). Since $G$ is generated by $2$-torsion elements this implies that $\pi_{1}(W,x)$ is trivial, or in other words that $W$ is simply connected. Since $W$ is a (ramified) covering of the non-compact Riemann surface ${\mathbb{H}}/\Gamma^{+}_{0}(\delta)$, $W$ cannot be compact. Hence $W$ is contractible by the uniformization theorem. This finishes the proof (2).
By Lemma 5.6 we now have
$\pi_{1}(V,x)=\ast_{i}(\langle{{\tau}}_{i}\rangle/({{\tau}}_{i}^{e_{i}}))$
Consequently for each $i$ we have a surjection
$G\rightarrow\langle{{\tau}}_{i}\rangle/({{\tau}}^{e_{i}}_{i})$
whose image lies in the $2$-torsion part of $\langle{{\tau}}_{i}\rangle/({{\tau}}_{i}^{e_{i}})$. It follows that $e_{i}=2$ for all $i$. We now apply Lemma 5.2 with $R=\{({{{\tau}}}_{i})_{i}\}\subset G$ and $S$ as in the statement of Lemma 5.8. From Lemma 5.2(2) we obtain (4). From Lemma 5.2(1) we obtain that $S$ is precisely the set of order two elements in $G$. We then get (3) from the definition of $S$.
To prove (1) we observe that by definition the fixed points of the elements of $G$ are orbits of the fixed points of the $({{\tau}}_{i})_{i}$. By Lemma 5.2(1) we have $\{({{\tau}}_{i})_{i}\}\subset S$. Hence the fixed points of $G$ are the fixed points of the elements of $S$ (since $S$ is closed under conjugation) which in turn are the orbits of the fixed points of the $(s_{i})_{i}$. This proves (1). ∎
Remark 5.9.
One may now give a topological realisation of the exact sequence (5.1). Let the notations be as in the statement and the proof of Lemma 5.8. Let $\mathcal{S}$$\subset{\mathbb{H}}$ be the set of fixed points of elements of $G$. By Lemma 5.8(1) we see that sending $s\in S$ to its unique fixed point provides a bijection between $S$ and $\mathcal{S}$}. By Lemma 5.8(1) we have a ramified $G$-covering ${\mathbb{H}}\rightarrow{\mathbb{H}}/G$ which restricts to an unramified $G$-covering ${\mathbb{H}}-{\mathcal{S}}\rightarrow V^{\circ}$ where as before $V^{\circ}$ is obtained from $V=[{\mathbb{H}}/G]$ by removing the orbifold points. We get a short exact sequence
(5.4) $1\rightarrow\pi_{1}({\mathbb{H}}-{\mathcal{S}},y)\xrightarrow{\varphi}\pi_{1}(V^{\circ},x)\rightarrow G\rightarrow 1$
If $\gamma$ is a path from $y$ to a point $y^{\prime}$ in $\mathcal{S}$} and $\bar{\gamma}$ is the image of $\gamma$ in $V^{\circ}$ then since $y^{\prime}$ is a ramification point of order two we have
(5.5) $\varphi(\tau_{\gamma})=\tau^{2}_{\gamma}.$
By Lemma 5.8(3) we have that ${\mathbb{H}}/G$ is contractible. Hence $\pi_{1}(V^{\circ},x)$ is the free group generated by $(\tau_{i})_{i}$, so its free basis is indexed by $R$ as in Lemma 5.3. On the other hand $\pi_{1}({\mathbb{H}}-{\mathcal{S}},y)$ is freely generated by a system of loops in ${\mathbb{H}}-{\mathcal{S}}$ associated to a system of non-intersecting paths from $y$ to the points in $\mathcal{S}$}, so its free basis is indexed by $S$ as in Lemma 5.3.
5.4. Application
Assume now we are in the standard setting of §4 where we have identified ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}\cong{\mathbb{H}}$, $\operatorname{SO}^{+}(q,{\mathbb{Z}})\cong\Gamma^{+}_{0}(\delta)$ and let $G$ be the group generated by reflections at roots in ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$. Then we have a presentation of $\Gamma^{+}_{0}(\delta)$ as in (5.3).
Let $(x_{i})_{i\in T}\subset(x_{i})_{i}$ be those fixed points (when lifted to $\mathbb{H}$}) which correspond to roots. Then $G$ is generated by the conjugates of $(\tau^{e_{i}/2}_{i})_{i\in T}$ ($e_{i}$ must be even since ${\mathbb{Z}}/2{\mathbb{Z}}$ is contained in the stabiliser).
Lemma 5.10.
Let $(G_{i})_{i}$ be a collection of groups with normal subgroups $H_{i}\subset G_{i}$. The kernel of $\ast_{i}G_{i}\rightarrow\ast_{i}(G_{i}/H_{i})$ is generated by the conjugates of the $H_{i}$.
Proof.
Let $K$ be the subgroup of $\ast_{i}G_{i}$ generated by the conjugates of $H_{i}$. Then $K$ is a normal subgroup and one checks that $(\ast_{i}G_{i})/K$ is the coproduct of the $(G_{i}/H_{i})_{i}$, by verifying the universal property. ∎
By Lemma 5.10 and the discussion preceding it, together with (5.3), we find that $G$ can be described as the kernel of $\Gamma^{+}_{0}(\delta)\twoheadrightarrow H$ where $H$ is the free product of the following factors
  1. (i)
    $\langle a_{i}\rangle$, $\langle b_{i}\rangle$ for $i=1,\ldots g$, $\langle\tau_{i}\rangle$ for $i=m+1,\ldots,m+n-1$.
  2. (ii)
    $\langle\tau_{i}\rangle/(\tau_{i}^{e_{i}})$ for $i=1,\ldots,m$, $i\not\in T$.
  3. (iii)
    $\langle\tau_{i}\rangle/(\tau_{i}^{e_{i}/2})$ for $i=1,\ldots,m$, $i\in T$.
Corollary 5.11.
$G$ is finitely generated if and only if all of the following conditions hold:
  1. (1)
    the genus of $X^{+}_{0}(\delta):={\mathbb{H}}/\Gamma^{+}_{0}(\delta)$ is zero;
  2. (2)
    $X^{+}_{0}(\delta)$ has one cusp;
  3. (3)
    there is at most one point $x\in X^{+}_{0}(\delta)$ with one of the following properties:
    1. (a)
      $x$ is a fixed point which is not a root point;
    2. (b)
      $x$ is a root point of order >2>2.
Proof.
According to [28, Theorem 1] $G$ will be finitely generated if and only if $H$ is finite, i.e. if it consists of a single group of type (ii) or (iii). This is equivalent to the description in the corollary. ∎
The description of $G$ as a kernel of a map between free products allows us to give explicit generators for $G$ via Lemma 5.12 below which is a more precise version of Lemma 5.10.
Lemma 5.12.
Let $A_{i}\subset G_{i}$ be right coset representatives for $H_{i}$. Let $G_{i}/H_{i}\rightarrow G_{i}:aH_{i}\mapsto G$ for $a\in A_{i}$ be the corresponding section and extend this to a section $\phi:\ast_{i}(G_{i}/H_{i})\rightarrow\ast_{i}G_{i}$. The kernel of $\ast_{i}G_{i}\rightarrow\ast_{i}(G_{i}/H_{i})$ is generated by $(\phi(u)H_{i}\phi(u)^{-1})_{i}$ for $u\in\ast_{i}(G_{i}/H_{i})$.
Proof.
One checks that the subgroup of $\ast_{i}G_{i}$ generated by $(\phi(u)H_{i}\phi(u)^{-1})_{i}$ is closed under by conjugation by $G_{i}$. ∎
6. Groups generated by spherical twists in $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$
6.1. Introduction
We start with some introductory comments. Throughout $Y$ is a K3 surface of Picard rank $1$ and degree $d=2\delta$. First observe the following.
Lemma 6.1.
[…]
are smooth orbifolds.
Proof.
By Lemma B.5 it suffices to prove this result for $\operatorname{Aut}^{+}{\mathcal{N}}(Y)$ instead of $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$ and by Corollary 4.3 combined with Remark A.6 we may further replace $({\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0},\operatorname{Aut}^{+}{\mathcal{N}}(Y))$ by $({\mathbb{H}},\Gamma^{+}_{0}(\delta))$. We conclude by the fact that $\Gamma^{+}_{0}(\delta)$ is a Fuchsian group (see §3.4). ∎
Lemma 6.2 (see also [19]).
Assume $\mathcal{G}$} is a subgroup of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ containing $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$ and let $G$ be its image in $\operatorname{Aut}^{+}H^{\ast}(Y,{\mathbb{Z}})$. We have an isomorphism
${\mathcal{G}}=\pi_{1}([{\mathcal{P}}_{0}^{+}(Y)/G])$
such that the long exact sequence of homotopy groups for the map ${\mathcal{P}}_{0}^{+}(Y)\rightarrow[{\mathcal{P}}_{0}^{+}(Y)/G]$} yields
(6.1) $1\rightarrow\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)\rightarrow{\mathcal{G}}\rightarrow G\rightarrow 1$
Proof.
Since $\mathcal{G}$} acts on the contractible space $\operatorname{Stab}^{\dagger}(Y)$ (see Theorem 3.20) we obtain from Theorem 3.20 a tautological isomorphism
${\mathcal{G}}=\pi_{1}([\operatorname{Stab}^{\dagger}(Y)/{\mathcal{G}}])=\pi_{1}([{\mathcal{P}}_{0}^{+}(Y)/G])$
and it follows from the general theory that this leads to the exact sequence (6.1). ∎
The following was also observed in [19] (see (1.2) on p.4):
Lemma 6.3.
Let $\mathcal{G}$} be a subgroup of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/N$ that contains $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$ and let $G$ be the image of $\mathcal{G}$} in $F(\delta)$ under the rightmost map in (B.4). Then
${\mathcal{G}}=\pi_{1}([{\mathbb{H}}^{0}/G])$
in such a way that the long exact sequence for homotopy groups for the map ${\mathbb{H}}^{0}\rightarrow[{\mathbb{H}}^{0}/G]$} is precisely
(6.2) $1\rightarrow\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]\rightarrow{\mathcal{G}}\rightarrow G\rightarrow 1$
Proof.
Put
$\operatorname{Stab}^{n}(Y):=\operatorname{Stab}^{\dagger}(Y)/\widetilde{\operatorname{GL}}_{2}^{+}({\mathbb{R}})$
[…]
Recall that ${\mathcal{P}}^{+}_{0}(Y)/\operatorname{GL}_{2}^{+}({\mathbb{R}})={\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}-\overline{\Delta^{+}(Y)}\cong{\mathbb{H}}^{0}$. We have a commutative diagram where the horizontal maps are coverings with the indicated Galois group.
$\operatorname{Stab}^{\dagger}(Y)$${\mathcal{P}}^{+}_{0}(Y)$$\operatorname{Stab}^{n}(Y)$${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}-\overline{\Delta^{+}(Y)}$$\scriptstyle{\widetilde{\operatorname{GL}}^{+}_{2}({\mathbb{R}})}$$\scriptstyle{\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)}$$\scriptstyle{\operatorname{GL}_{2}^{+}({\mathbb{R}})}$$\scriptstyle{\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]}$
Then we find in a similar way as in Lemma 6.2
${\mathcal{G}}=\pi_{1}([{\mathbb{H}}^{0}/G])$
and this isomorphism induces (6.2). ∎
We will use the following result.
Lemma 6.4.
Let $Y$ be a K3 surface with Picard rank 1. Let ${\mathcal{G}}=\langle T_{S}\mid S\in\operatorname{Sph}(Y)\rangle\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ the subgroup generated by spherical twists and let $G$ be its image in $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$. Let $(S_{i})_{i\in I}\in\operatorname{Sph}(Y)$ be such that $\{v(S_{i})_{i}\}$ intersects all $G$-orbits in $\Delta(Y)$. Then we have:
  1. (1)
    if $(T_{S_{i}})_{i}$ generates the image $\bar{{\mathcal{G}}}$ of $\mathcal{G}$} in $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/N$ then $(T_{S_{i}})_{i}$ generates $\mathcal{G}$};
  2. (2)
    if in addition the $T_{S_{i}}$ yield free generators for $\bar{{\mathcal{G}}}$ then they are free generators for $\mathcal{G}$}.
Proof.
The claim (2) is an obvious consequence of (1), so we concentrate on the latter. The hypotheses imply that $\mathcal{G}$} is generated by $(T_{S_{i}})_{i}$ and $N$. In other words if $S$ is a spherical object in ${\mathcal{D}}^{b}(Y)$ then by Corollary 3.23 and Lemma B.2(3) $S$ can be written as
$S\cong T_{S_{i_{1}}}^{\pm}\cdots T_{S_{i_{n}}}^{\pm}S_{j}[2n]$
for a suitable $S_{j}$ and $n\in{\mathbb{Z}}$. But then by Lemma 3.10 we have
$T_{S}=T_{S[-2n]}=T_{S_{i_{1}}}^{\pm 1}\cdots T_{S_{i_{n}}}^{\pm 1}T_{S_{j}}T_{S_{i_{n}}}^{\mp 1}\cdots T_{S_{i_{1}}}^{\mp 1}$
In other words $T_{S}$ is contained in the subgroup of $\mathcal{G}$} generated by $(T_{S_{i}})_{i}$. Since this is true for any $T_{S}$, this subgroup must in fact be equal to $\mathcal{G}$}. ∎
6.2. Main result
Theorem 6.5.
Let $Y$ be a K3 surface of Picard rank $1$ and degree $d=2\delta$. Let ${\mathcal{G}}=\langle T_{S}\ |S\in\operatorname{Sph}(Y)\rangle\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$} be the group generated by all spherical twists and let $G\subset\operatorname{Aut}^{+}{\mathcal{N}}(Y)/\{\pm 1\}=\Gamma^{+}_{0}(n)$ be the corresponding group generated by all spherical reflections. Then $\mathcal{G}$} is freely generated by
$\langle T_{E_{i}}\ |\ v(E_{i})=\rho_{i}\in\Delta^{+}(Y);i\in I\rangle,$
where $E_{i}$ is the vector bundle with $v(E_{i})=\rho_{i}$ and $(\rho_{i})_{i}$ is a choice of orbit representatives, up to sign, for $G$ acting on the roots, such that the set $\{x_{i}=p(\bar{\rho_{i}})\}_{i\in I}\in{\mathbb{H}}$ (see (4.7) for the definition of $p$) is contained in a convex fundamental domain $\mathcal{F}$} for $G$, containing $\infty$ in its closure in $\overline{{\mathbb{H}}}$ (note that such a fundamental domain exists by Lemma 3.28 and the fact that $G$ is a normal subgroup).
Proof.
To simplify the discussion we will identify ${\mathbb{P}}({\mathcal{N}}(Y)_{\mathbb{R}})_{<0}$ and $\mathbb{H}$}. E.g. when we speak about “roots” in $\mathbb{H}$} we mean points of the form $p(\bar{\rho})$ where $\rho\in\Delta^{+}(Y)$. As usual ${\mathbb{H}}^{0}$ is the complement of ${\mathcal{S}}:=\{p(\bar{\rho})\mid\rho\in\Delta^{+}(Y)\}$. Recall that by Lemma 5.8(1), the points of $\mathcal{S}$} are order two ramification points for ${\mathbb{H}}\rightarrow{\mathbb{H}}/G$ and furthermore $G$ acts freely on ${\mathbb{H}}^{0}$ so ${\mathbb{H}}^{0}/G]={\mathbb{H}}^{0}/G$. By Lemma it is sufficient to prove the theorem with $\mathcal{G}$} replaced by $\bar{{\mathcal{G}}}$, the image of ${\mathcal{G}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. By Theorem 3.21 $\bar{{\mathcal{G}}}$ contains $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$. By Lemma 6.3 the exact sequence
(6.3) $1\rightarrow\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]\rightarrow\bar{{\mathcal{G}}}\rightarrow G\rightarrow 1$
is obtained from the long exact sequence of homotopy groups associated with ${\mathbb{H}}^{0}\rightarrow{\mathbb{H}}^{0}/G$}. I.e. after choosing a base point $y\in{\mathbb{H}}^{0}$ with image $x\in{\mathbb{H}}^{0}/G$, (6.3) may be identified with
(6.4) $1\rightarrow\pi_{1}({\mathbb{H}}^{0},y)\xrightarrow{\varphi}\pi_{1}({\mathbb{H}}^{0}/G,x)\rightarrow G\rightarrow 1$
Let $\{\rho_{i}\}_{i\in I}\in{\mathbb{H}}$ be orbit representatives for the roots contained in $\mathcal{F}$}. Note that all $\rho_{i}$ are automatically on the boundary of $\mathcal{F}$} and the hypothesis on $\mathcal{F}$} imply that the interior of $\mathcal{F}$} does intersect the vertical line segments discussed in Remark 3.22. Choose the base point $y$ in the interior of $\mathcal{F}$}. Then as explained in §5.3, $\pi_{1}({\mathbb{H}}^{0}/G,x)$ is the free group freely generated by loops $(\tau_{\overline{\gamma}_{i}})_{i}$ around the punctures corresponding to the images $(\overline{\gamma}_{i})_{i}$ of a system of non-intersecting paths $(\gamma_{i})_{i}$ in $\mathcal{F}$} connecting $y$ to $\rho_{i}$. Recall from (5.5) that we have $\varphi(\tau_{\gamma_{i}})=\tau^{2}_{\overline{\gamma}_{i}}$. We will now assume that the paths $\gamma_{i}$ have been chosen as in Remark 3.22. Then under the identification of (6.3) and (6.4) $\tau_{\gamma_{i}}$ corresponds to $T_{E_{i}}^{2}$ considered as an element of $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]$ (see Theorem 3.21). It then follows that $\tau^{2}_{\overline{\gamma}_{i}}$ corresponds to $T_{E_{i}}^{2}$ considered as an element of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/N$. Since square roots are unique in free groups we obtain that $\tau_{\overline{\gamma}_{i}}$ corresponds to $T_{E_{i}}$. ∎
7. Examples
In this section we explicitly describe the subgroup ${\mathcal{G}}\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ generated by spherical twists for some specific examples of K3 surfaces with Picard rank $1$. The first four examples we consider are general anticanonical divisors $Y$ in Fano threefolds $X$ of Picard rank $1$ possessing full exceptional collections of vector bundles, namely, $X$ is one of the four types ${\mathbb{P}}^{3},Q_{3},V_{5}$ and $V_{22}$ (see Lemma 3.5). By a theorem due to Moishezon [38, Theorem 7.5] (see also [54, Théorème 15.33]) such a K3 surface also has Picard rank $1$. In each of these cases we show that $\mathcal{G}$} is the free group $F_{4}$ generated by $T_{F_{0}|_{Y}},\dots,T_{F_{3}|_{Y}}$ for a fixed exceptional collection ${\mathcal{F}}=(F_{0},\dots,F_{3})$ on the ambient Fano threefold $X$. A similar result is established for the case of a K3 surface of degree $2$ arising as a double cover of ${\mathbb{P}}^{2}$. Finally, we give an example of a K3 surface (of degree 8) for which the group $\mathcal{G}$} is not finitely generated but we still list explicit generators.
By Theorem 6.5 to find a set of free generators of $\mathcal{G}$} in each of the cases it is sufficient to find a set of orbit representatives for the fixed points in $\mathbb{H}$} of elements of $G$ contained in an appropriate fundamental domain, where $G\subset\operatorname{Aut}^{+}\operatorname{H}^{*}(Y)$ is the group generated by spherical reflections. The fundamental domains in this section are precisely the “Ford domains” which were defined in the proof of Lemma 3.28. We obtain such a fundamental domain explicitly from a fundamental domain for the corresponding group $\Gamma_{0}^{+}(\delta)$ using Lemma 3.29. In turn, since in all of our cases $\Gamma_{0}^{+}(\delta)=F(\delta)$ is the Fricke group, constructing a fundamental domain for $\Gamma_{0}^{+}(\delta)$ amounts to finding a fundamental domain for $\Gamma_{0}(\delta)$ invariant under the Fricke involution and cutting it along the half-circle with radius $1/\sqrt{\delta}$ centred at 0. These fundamental domains for $\Gamma_{0}^{+}(\delta)$ can be also found in e.g. [5], [51].
Below we give a brief summary of the results and then discuss each example in detail.
Theorem 7.1.
  1. (1)
    Let $X$ be a Fano threefold with Picard rank $1$ and $h^{2,1}=0$, that is, $X={\mathbb{P}}^{3},Q_{3},V_{5}$} or $V_{22}$ (the constructions are reviewed below). Let $Y\subset X$ be a generic anticanonical divisor which is a K3 surface of Picard rank $1$ and degree $4,6,10$ and $22$ respectively. Then the group
    ${\mathcal{G}}=\langle T_{S}\ |\ v(S)\in\Delta(Y)\rangle\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$
    generated by all spherical twists is isomorphic to the free group $F_{4}$ on $4$ generators. In each case we can give an explicit set of free generators which are the spherical twists along the spherical vector bundles obtained by restricting a full exceptional collection on vector bundles on $X$ to $Y$.
  2. (2)
    Let $Y$ be a K3 surface of degree $2$ which is a double cover of ${\mathbb{P}}^{2}$ branched along a sextic curve. Then the group ${\mathcal{G}}\cong F_{3}$ is freely generated by the spherical twists along the pullbacks of $3$ lines bundles forming a full exceptional collection on ${\mathbb{P}}^{2}$.
  3. (3)
    Let $X$ be a complete intersection of $3$ quadrics in ${\mathbb{P}}^{6}$ and $Y\subset X$ a smooth anticanonical section. $Y$ is a K3 surface of degree $8$. Then $\mathcal{G}$} is freely generated by $T_{{\mathcal{O}}_{Y}(k)}$ for all $k\in{\mathbb{Z}}$ (in particular, it is infinitely generated).
7.1. $Y\subset{\mathbb{P}}^{3}$ (degree 4)
Let $X={\mathbb{P}}^{3}$. Then a smooth anticanonical divisor $Y\subset X$ is a quartic K3 surface. A standard full exceptional collection on $X$ consisting of line bundles is due to Beilinson [6]:
${\mathcal{F}}=({\mathcal{O}},{\mathcal{O}}(1),{\mathcal{O}}(2),{\mathcal{O}}(3))$
The quadratic form corresponding to the Mukai pairing on ${\mathcal{N}}(Y)$ is $q(r,l,s)=2l^{2}-rs$. The spherical vector bundles ${\mathcal{O}}(i)|_{Y}$ on $Y$ have the following classes in ${\mathcal{N}}(Y)$:
$v({\mathcal{O}}|_{Y})=(1,0,1),\ v({\mathcal{O}}(1)|_{Y})=(1,1,3),\ v({\mathcal{O}}(2)|_{Y})=(1,2,9),\ v({\mathcal{O}}(3)|_{Y})=(1,3,19)$
Using the formula in Corollary 4.4, one finds that the corresponding $4$ points in $\mathbb{H}$} are
$\frac{i}{\sqrt{2}},1+\frac{i}{\sqrt{2}},2+\frac{i}{\sqrt{2}},3+\frac{i}{\sqrt{2}}$
The group $\Gamma_{0}^{+}(2)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has one elliptic point of order $2$, one elliptic point of order $4$ and one cusp ([16, Table 4]). $X_{0}^{+}(2)$ has genus 0. The elliptic elements of $\Gamma_{0}^{+}(2)$ can be understood in the following way. Let $s_{0}\in\Gamma_{0}^{+}(2)$ be the transformation corresponding to $T_{{\mathcal{O}}_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$, namely $s_{0}=\begin{pmatrix}0&-1\\ 2&0\end{pmatrix}$, $t:z\mapsto\frac{-1}{2z}$. The element $s_{0}$ generates an elliptic subgroup of order $2$ in $\Gamma_{0}^{+}(2)$. Its fixed point in $\mathbb{H}$} is $\frac{i}{\sqrt{2}}$. Now let $t\in\Gamma_{0}^{+}(2)$ be the transformation $t=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$, $t:z\mapsto z+1$. Observe that by Corollary 4.6 $t$ corresponds to the autoequivalence
$\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y).$
We claim that $s_{0}t$ generates an elliptic subgroup of order $4$. Indeed, note that $s_{k}:=t^{k}s_{0}t^{-k}$ corresponds to the spherical twist $T_{{\mathcal{O}}_{Y}(k)}$. Hence $(s_{0}t)^{4}=s_{0}s_{1}s_{2}s_{3}t^{4}=1$ by Proposition 3.13, since $\omega_{X}\cong{\mathcal{O}}_{X}(-4)$. Note that up to even shifts this relation holds not only numerically, but on the level of derived categories as well. By the discussion after Lemma 5.10 the subgroup $G\subseteq\Gamma_{0}^{+}(2)$ generated by reflections is the kernel
$1\to G\to\Gamma_{0}^{+}(2)\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/4{\mathbb{Z}}\xrightarrow{s_{0}\mapsto 0,t^{4}\mapsto 0}{\mathbb{Z}}/4{\mathbb{Z}}\to 1$
In particular, $\Gamma_{0}^{+}(2):G]=4$ and it is easy to see that $G$ is the free product of $4$ copies of ${\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0},ts_{0}t^{-1},t^{2}s_{0}t^{-2}$ and $t^{3}s_{0}t^{-3}$. These elements correspond to the spherical twists with respect to exactly the restrictions of the vector bundles of the exceptional collection $\mathcal{F}$}. Figure 1 depicts a fundamental domain for $G$ which contains the corresponding $4$ fixed points of these generators.
Figure 1. A fundamental domain for $G\subset\Gamma_{0}^{+}(2)$ containing $\frac{i}{\sqrt{2}},1+\frac{i}{\sqrt{2}},2+\frac{i}{\sqrt{2}}$} and $3+\frac{i}{\sqrt{2}}$.
7.2. $Y\subset Q_{3}$ (degree 6)
Let $X$ be a smooth quadric threefold. Then a smooth anticanonical divisor $Y\subset X$ is a K3 surface of degree $6$. Kapranov [27] constructed the following full exceptional collection of vector bundles on $Q_{3}$:
${\mathcal{F}}=({\mathcal{O}},{\mathcal{S}}^{*},{\mathcal{O}}(1),{\mathcal{O}}(2)),$
where $\mathcal{S}$} is the spinor bundle. The quadratic form corresponding to the Mukai pairing on ${\mathcal{N}}(Y)$ is $q(r,l,s)=3l^{2}-rs$. The spherical vector bundles on $Y$ corresponding to the objects of $\mathcal{F}$} have the following classes in ${\mathcal{N}}(Y)$:
$v({\mathcal{O}}|_{Y})=(1,0,1),\ v({\mathcal{S}}^{*}|_{Y})=(2,1,2),\ v({\mathcal{O}}(1)|_{Y})=(1,1,4),\ v({\mathcal{O}}(2)|_{Y})=(1,2,13)$
Using the formula in Corollary 4.4, one finds that the corresponding $4$ points in $\mathbb{H}$} are
(7.1) $\frac{i}{\sqrt{3}},\frac{1}{2}+\frac{i}{2\sqrt{3}},1+\frac{i}{\sqrt{3}},2+\frac{i}{\sqrt{3}}$
The group $\Gamma_{0}^{+}(3)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has one elliptic point of order $2$, one elliptic point of order $6$ and one cusp ([16], Table 4). $X_{0}^{+}(3)$ has genus 0. Again, let $s_{0}\in\Gamma_{0}^{+}(3)$ be the transformation corresponding to $T_{{\mathcal{O}}_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ and $t\in\Gamma_{0}^{+}(3)$ corresponding to $\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Then, similarly to the previous example, by Proposition 3.13 we have $(ts_{0})^{3}=s$, where $s$ denotes the transformation corresponding to $T_{{\mathcal{S}}^{*}|_{Y}}$. Hence $ts_{0}$ generates an elliptic subgroup of order $6$. Same as in the previous example, up to an even shift, the autoequivalence ${\mathcal{O}}_{Y}(1)\otimes T_{{\mathcal{O}}_{Y}}(-)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ has order $6$ as well. The subgroup $G$ generated by reflections has index $3$ in $\Gamma_{0}^{+}(3)$:
$1\to G\to\Gamma_{0}^{+}(3)\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/6{\mathbb{Z}}\xrightarrow{s_{0}\mapsto 0,t^{3}\mapsto 0}{\mathbb{Z}}/3{\mathbb{Z}}\to 1$
$G\cong({\mathbb{Z}}/2{\mathbb{Z}})^{\ast 4}$ is generated by $s_{0},ts_{0}t^{-1},t^{2}s_{0}t^{-2},(ts_{0})^{3}$ (we are listing the generators corresponding to the chosen roots). Figure 2 depicts a fundamental domain of $G$ which contains the points (7.1). It is constructed by the same procedure as in the case of $Y\subset{\mathbb{P}}^{3}$.
Figure 2. A fundamental domain for $G\subset\Gamma_{0}^{+}(3)$ containing $\frac{i}{\sqrt{3}},\frac{1}{2}+\frac{i}{2\sqrt{3}},1+\frac{i}{\sqrt{3}}$ and $2+\frac{i}{\sqrt{3}}$.
7.3. $Y\subset V_{5}$ (degree 10)
Let $X$ be a section of the Plücker embedding of $\operatorname{Gr}(2,5)$ by a subspace of codimension $3$. $X$ is a Fano variety of type $V_{5}$ and a smooth anticanonical divisor $Y\subset X$ is a K3 surface of degree $10$. We consider a full exceptional collection on $X$ constructed by Orlov in [42], namely,
${\mathcal{F}}=({\mathcal{O}},{\mathcal{Q}}_{X},{\mathcal{U}}^{*}_{X},{\mathcal{O}}(1)),$
where $\mathcal{U}$} and $\mathcal{Q}$} are the universal and the quotient bundles on $\operatorname{Gr}(2,5)$ respectively. Applying the mutation $\sigma_{1}$ to $\mathcal{F}$} one obtains an exceptional collection of shifted vector bundles whose restrictions to $Y$ have the following classes in ${\mathcal{N}}(Y)$:
(7.2) $(-2,1,-3),(1,0,1),(2,1,3),(1,1,6).$
Since shifting a spherical object does not affect the corresponding spherical twist, we can replace the first object $F_{0}^{\prime}$ by $F_{0}^{\prime}[1]$. This amounts to changing the sign of the first vector, making all ranks positive. And the $4$ points in $\mathbb{H}$} are
(7.3) $\frac{-1}{2}+\frac{i}{2\sqrt{5}},\frac{i}{\sqrt{5}},\frac{1}{2}+\frac{i}{2\sqrt{5}},1+\frac{i}{\sqrt{5}}.$
The group $\Gamma_{0}^{+}(5)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has three elliptic points of order $2$ and one cusp ([16, Table 4]). $X_{0}^{+}(5)$ has genus 0. The elements $s_{0},s_{1}\in\Gamma_{0}^{+}(5)$ corresponding to $T_{{\mathcal{O}}_{Y}}$ and $T_{{\mathcal{Q}}|_{Y}}$ respectively generate two non-congruent elliptic subgroups of order $2$. The remaining elliptic subgroup of order $2$ is generated by $s_{0}s_{1}t\in\Gamma_{0}^{+}(5)$, where as usual $t$ corresponds to ${\mathcal{O}}_{Y}(1)\otimes(-)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. This element is not a point reflection with respect to a root. One can verify that $(s_{0}s_{1}t)^{2}=1$ holds on the level of derived categories as well, namely, up to an even shift the autoequivalence $\psi(-):=T_{{\mathcal{O}}_{Y}}T_{{\mathcal{Q}}|_{Y}}(-\otimes{\mathcal{O}}_{Y}(1))$ has order two. Indeed, it is easy to check that up to a shift ${\mathcal{U}}^{*}|_{Y}={\mathcal{O}}_{Y}(1)\otimes T_{{\mathcal{O}}_{Y}}({\mathcal{Q}}|_{Y})$. Using this expression to write $T_{{\mathcal{U}}^{*}}$ in terms of $T_{{\mathcal{O}}_{Y}}$, $T_{{\mathcal{Q}}|_{Y}}$ and ${\mathcal{O}}_{Y}(1)\otimes(-)$ and substituting to the formula in Proposition 3.13 with $\omega_{X}={\mathcal{O}}_{X}(-2)$, we get that $\psi$ has order $2$ up to an even shift. By the discussion after Lemma 5.10 the subgroup $G$ generated by reflections is the kernel
$1\to G\to\Gamma_{0}^{+}(5)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{\ast 3}\xrightarrow{s_{0}\mapsto 0,s_{1}\mapsto 0,t^{2}\mapsto 0}{\mathbb{Z}}/2{\mathbb{Z}}\to 1.$
One can see that $G\cong({\mathbb{Z}}/2{\mathbb{Z}})^{\ast 4}$ is generated by $s_{0},s_{1},ts_{0}t^{-1}$ and $ts_{1}t^{-1}$ (the generators corresponding to the roots (7.2) are $s_{0},s_{0}s_{1}s_{0},ts_{0}t^{-1}$ and $ts_{0}s_{1}s_{0}t^{-1}$). Figure 3 shows a fundamental domain for $G$ containing the four points (7.3).
Figure 3. A fundamental domain for $G\subseteq\Gamma_{0}^{+}(5)$ containing $\frac{-1}{2}+\frac{i}{2\sqrt{5}},\frac{i}{\sqrt{5}},\frac{1}{2}+\frac{i}{2\sqrt{5}}$ and $1+\frac{i}{\sqrt{5}}$.
7.4. $Y\subset V_{22}$ (degree 22)
Let $X$ be the zero locus of $(\Lambda^{2}{\mathcal{U}}^{*})^{\oplus 3}$ in $\operatorname{Gr}(3,7)$, where $\mathcal{U}$} denotes the universal bundle. $X$ a Fano variety of type $V_{22}$ and a smooth anticanonical divisor $Y\subset X$ is a K3 surface of degree $22$. A full exceptional collection of vector bundles on $X$ was constructed by Kuznetsov in [33] (see [34, Theorem 2] for the proof of the fullness, also see [35, Theorem 4.1]):
${\mathcal{F}}=({\mathcal{O}},{\mathcal{U}}^{*}_{X},E^{*},\Lambda^{2}{\mathcal{U}}^{*}_{X}).$
Here $E$ is a vector bundle of rank $2$, see e.g. [33] for the construction. Applying a sequence of $3$ mutations, we obtain another exceptional collection, which automatically consists of shifted vector bundles by Proposition 3.4:
${\mathcal{F}}^{\prime}=\sigma_{1}\sigma_{2}\sigma_{3}{\mathcal{F}}=(L_{{\mathcal{O}}}L_{{\mathcal{U}}^{*}_{X}}L_{E^{*}}\Lambda^{2}{\mathcal{U}}^{*}_{X},{\mathcal{O}},{\mathcal{U}}^{*}_{X},E^{*}).$
One can check that the corresponding spherical objects on $Y$ have the following classes in ${\mathcal{N}}(Y)$:
(3,1,4),(1,0,1),(3,1,4),(2,1,6)(-3,1,-4),\,(1,0,1),(3,1,4),(2,1,6)
As in the previous example, we are free to change the sign of the first vector, making all ranks positive. The corresponding $4$ points in $\mathbb{H}$} are
$\frac{-1}{3}+\frac{i}{3\sqrt{11}},\frac{i}{\sqrt{11}},\frac{1}{3}+\frac{i}{3\sqrt{11}},\frac{1}{2}+\frac{i}{2\sqrt{11}}.$
The group $\Gamma_{0}^{+}(11)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has $4$ elliptic points of order $2$ and one cusp ([16], Table 4). $X_{0}^{+}(11)$ has genus 0.
Figure 4. A fundamental domain for $G=\Gamma_{0}^{+}(11)$ containing $\frac{-1}{3}+\frac{i}{3\sqrt{11}},\frac{i}{\sqrt{11}},\frac{1}{3}+\frac{i}{3\sqrt{11}},$ and $\frac{1}{2}+\frac{i}{2\sqrt{11}}$.
7.5. Double cover of ${\mathbb{P}}^{2}$ (degree 2)
Let $\pi:Y\to{\mathbb{P}}^{2}$ be a double covering branched along a sextic curve. Then $Y$ is a K3 surface of degree $2$. Pulling back the standard exceptional collection $({\mathcal{O}},{\mathcal{O}}(1),{\mathcal{O}}(2))$ from ${\mathbb{P}}^{2}$ to $Y$, we obtain a collection of $3$ spherical vector bundles with the following classes in ${\mathcal{N}}(Y)$:
$(1,0,1),(1,1,2),(1,2,5)$
[…]
By the discussion after Lemma 5.10 the group $G$ generated by reflections is the kernel
$1\to G\to\Gamma_{0}^{+}(1)\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/3{\mathbb{Z}}\xrightarrow{s_{0}\mapsto 0,t^{3}\mapsto 0}{\mathbb{Z}}/3{\mathbb{Z}}\to 1$
One observes that $\Gamma_{0}^{+}(1):G]=3$ and $G\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0}$, $ts_{0}t^{-1}$ and $t^{2}s_{0}t^{-2}$. These elements correspond to spherical twists with respect to the pullbacks of the bundles forming the standard exceptional collection on ${\mathbb{P}}^{2}$. Figure 5 shows a fundamental domain for $G$ containing the corresponding $3$ fixed points.
Figure 5. A fundamental domain for $G\subset\Gamma_{0}^{+}(1)\cong\operatorname{PSL}_{2}({\mathbb{Z}})$ containing $i,1+i$ and $2+i$.
7.6. An example of an infinitely generated group of spherical twists (degree 8)
Let $X$ be a complete intersection of $3$ quadrics in ${\mathbb{P}}^{6}$. Then $X$ is a Fano variety of Picard rank $1$, index $1$ and degree $8$. A smooth anticanonical section $Y\subset X$ is a K3 of degree $8$. The quadratic form corresponding the the Mukai pairing on ${\mathcal{N}}(Y)$ is $q(r,l,s)=4l^{2}-rs$. By [16, Table 4] the group $\Gamma_{0}^{+}(4)$ has one elliptic point of order $2$ and two cusps, so by Corollary 5.11 the subgroup $G\subset\Gamma_{0}^{+}(4)$ generated by reflections has infinite index and is infinitely generated. Hence the subgroup $\mathcal{G}$} of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)/[1]$ generated by all spherical twists is infinitely generated as well. by Theorem 6.5. As before, the element $s_{0}\in\Gamma_{0}^{+}(4)$ corresponding to $T_{{\mathcal{O}}_{Y}}$ generates an elliptic subgroup of order $2$. Its fixed point is $\frac{i}{2}$ and by the discussion after Lemma 5.10 the group $G=\ast_{n\in{\mathbb{Z}}}{\mathbb{Z}}/2{\mathbb{Z}}$ is the kernel
$1\to G\to{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}\xrightarrow{s_{0}\mapsto 1}{\mathbb{Z}}\to 1$
$G$ is generated by the conjugates $t^{k}s_{0}t^{-k}$, $k\in{\mathbb{Z}}$ corresponding to the spherical twists $T_{{\mathcal{O}}(k)_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Figure 6 shows a fundamental domain of $G$ containing the corresponding fixed points $\{k+\frac{i}{2}\}_{k\in{\mathbb{Z}}}$. We conclude that ${\mathcal{G}}\cong\langle T_{{\mathcal{O}}_{Y}(k)}\ |\ k\in{\mathbb{Z}}\rangle$ is freely generated by the spherical twists with respect to all line bundles.
Figure 6. A fundamental domain for $G\subset\Gamma_{0}^{+}(4)$ containing the fixed points $\{k+\frac{i}{2}\}_{k\in{\mathbb{Z}}}$
Remark 7.2.
In other examples considered in this section the associated Fano variety had a full exceptional collection. This is not the case for a complete intersection $X$ of $3$ quadrics in ${\mathbb{P}}^{6}$, since $h^{2,1}(X)=14\neq 0$ (see [26, §12.2]), so the Hodge diamond of $X$ is not diagonal. However, the derived category of $X$ is known to have a non-trivial semiorthogonal decomposition ${\mathcal{D}}^{b}(X)=\langle{\mathcal{O}}_{X},{\mathcal{D}}^{b}(\mathop{\text{{coh}}}{\mathcal{B}})\rangle$, where $\mathcal{B}$} is the even part of the sheaf of Clifford algebras ([8, Theorem 6.1]).
Remark 7.3.
The group $\mathcal{G}$} can be finitely generated even if the K3 surface in question is an anticanonical section of a Fano threefold without a full exceptional collection. For instance, let $X$ be a section of the Plücker embedding of $\operatorname{Gr}(2,6)$ by a subspace of codimension $5$. Then $X$ is a Fano threefold of Picard rank $1$, index $1$ and degree $14$ and a smooth anticanonical section $Y\subset X$ is a K3 surface of degree $14$. For such $Y$ the group $\mathcal{G}$} is finitely generated by Corollary 5.11, because $\Gamma_{0}^{+}(7)$ has genus 0, 22 elliptic points of order $2$, which are both root points, $1$ elliptic point of order $3$ and $1$ cusp ([16, Table 4]). On the other hand, $X$ has no full exceptional collection since its Hodge diamond is not diagonal ($h^{2,1}=5$, see [26, §12.2]).
8. Exceptional collections on Fano threefolds
Theorem 8.1 (Polishchuk, [44, Theorem 1.2]).
Let $X$ be a Fano threefold with Picard rank $1$ and very ample anticanonical class. Let $E,E^{\prime}$ be exceptional vector bundles on $X$ with the same restriction to a smooth anticanonical divisor $Y$. Assume that $\operatorname{Ext}^{1}(E,E\otimes\omega_{X}^{-1})=0$. Then $E\cong E^{\prime}$.
Remark 8.2.
Theorem 1.2 in the actual text [44] assumes that $E$ and $E^{\prime}$ have same class in $K_{0}(X)$ rather than the same restriction. However, examining Polishchuk’s proof one can see that the first step is to derive $E_{Y}=E^{\prime}_{Y}$ from this condition and afterwards the fact that $E$ and $E^{\prime}$ have the same class in $K_{0}(X)$ is not used.
Remark 8.3.
As observed in [44], the condition $\operatorname{Ext}^{1}(E,E\otimes\omega_{X}^{-1})=0$ is satisfied in particular when $E$ is included in a full exceptional collection of vector bundles, because on variety $X$ with $\operatorname{rk}K_{0}(X)=\dim X+1$ every full exceptional collection extends to a geometric helix ([11, Proposition 3.3]).
Lemma 8.4.
Let $X$ be a Fano variety and $i:Y\hookrightarrow X$ a smooth anticanonical divisor. Let $E$ be an exceptional object and $F$ an exceptional vector bundle in ${\mathcal{D}}^{b}(X)$. Assume $T_{i^{*}E}\cong T_{i^{*}F}$. Then $E$ is a shifted vector bundle.
Proof.
By Lemma 3.11 we have $i^{*}E\cong i^{*}F[n]$. By shifting $E$ we may assume without loss of generality assume that $n=0$. Having done this it remains to show that $E$ has one nonzero cohomology sheaf in degree 0. Consider the distinguished triangle
(8.1) $E(-Y)\to E\to i_{*}{\mathbb{L}}i^{*}E\to$
Suppose ${\mathcal{H}}^{l}(E)\neq 0$ for $l\neq 0$. We have ${\mathcal{H}}^{l}(i_{*}{\mathbb{L}}i^{*}E)=0$, hence the long exact sequence associated to (8.1) implies that the cokernel of ${\mathcal{H}}^{l}(E(-Y))={\mathcal{H}}^{l}(E)(-Y)\to{\mathcal{H}}^{l}(E)$} vanishes. On the other hand, this cokernel is isomorphic to ${\mathcal{H}}^{l}(E)|_{Y}$. Since $Y$ is an ample divisor this implies that ${\mathcal{H}}^{l}(E)$ must have finite support. Hence, it suffices to show that a non-zero cohomology sheaf of an exceptional object on $X$ cannot have zero-dimensional support. Suppose to the contrary that it does. For every point of $p\in X$ there exists a smooth anticanonical divisor passing through $p$, so we can pick another $i^{\prime}:Y^{\prime}\hookrightarrow X$ such that ${\mathcal{H}}^{l}(E)|_{Y^{\prime}}\neq 0$. By Nakayama ${\mathcal{H}}^{l}(E)(-Y^{\prime})\rightarrow{\mathcal{H}}^{l}(E)$ cannot be surjective. It follows that the spherical object $i^{\prime}_{*}{\mathbb{L}}i^{\prime*}E$ also has a cohomology sheaf of finite length. But the later is impossible, because cohomology sheaves of spherical objects are rigid (this is well known, see e.g. [25, Proposition 3.5]). ∎
Lemma 8.5.
Let $Y$ be a smooth variety and let ${\mathcal{G}}=\langle T_{S}\mid S\in\operatorname{Sph}(Y)\rangle\subset\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ be the subgroup generated by all spherical twists. Assume there are $(S_{i})_{i\in I}\in\operatorname{Sph}(Y)$ such that:
  1. (1)
    $\mathcal{G}$} is freely generated by $(T_{S_{i}})_{i}$;
  2. (2)
    for every $S\in\operatorname{Sph}(Y)$ there is some $H\in{\mathcal{G}}$, $i\in I$ and $n\in{\mathbb{Z}}$ such that $S\cong H(S_{i})[n]$.
Let
$\phi:=(T_{S^{(1)}},\dots,T_{S^{(n)}}),\qquad\phi^{\prime}:=(T_{S^{(1^{\prime})}},\dots,T_{S^{(n^{\prime})}})$
be $n$-tuples of spherical twists with the same product
(8.2) $T_{S^{(1)}}\cdots T_{S^{(n)}}=T_{S^{(1^{\prime})}}\cdots T_{S^{(n^{\prime})}}$
in $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Then $\phi$ and $\phi^{\prime}$ are in the same $B_{n}$-orbit under the Hurwitz action (see §3.2) on ${\mathcal{G}}^{n}$.
Proof.
Let ${\mathcal{G}}_{ab}={\mathcal{G}}/[{\mathcal{G}},{\mathcal{G}}]$ be the abelianisation of $G$. For $S\in\operatorname{Sph}(Y)$ we will denote by $\bar{S}$ the class of $T_{S}$ in ${\mathcal{G}}_{ab}$. Then ${\mathcal{G}}_{ab}$ is a free abelian group with a distinguished basis given by $(\bar{S}_{i})_{i\in I}$. By (2) and Lemma 3.10 for every $S\in\operatorname{Sph}(Y)$ there exists $H\in{\mathcal{G}}$ such that $T_{S}=HT_{S_{i}}H^{-1}$ and hence $\bar{S}=\bar{S}_{i}$ is an element of the distinguished basis. We also deduce from (2) that if $\bar{S}=\bar{S}^{\prime}$ then $T_{S}=JT_{S^{\prime}}J^{-1}$ for some $J\in{\mathcal{G}}$.
By (8.2) one has
$\bar{S}^{(1)}+\cdots+\bar{S}^{(n)}=\bar{S}^{(1^{\prime})}+\cdots+\bar{S}^{(n^{\prime})}$
Since all $\bar{S}^{(i)}$ and $\bar{S}^{(i^{\prime})}$ are in the distinguished basis, we have $\bar{S}^{(i)}=\bar{S}^{(\tau(i)^{\prime})}$ for some permutation $\tau$ of $\{1,\ldots,n\}$. As explained above $\bar{S}^{(i)}=\bar{S}^{(\tau(i)^{\prime})}$ implies $T_{S^{(i)}}=J_{i}T_{S^{(\tau(i)^{\prime})}}J_{i}^{-1}$ for some $J_{i}\in{\mathcal{G}}$ for every $i\in\{1,\dots,n\}$. It now follows from (1), (8.2) and Theorem 3.18 that $\phi$ and $\phi^{\prime}$ are in the same Hurwitz orbit. ∎
Theorem 8.6.
Let $X$ be a threefold with a full exceptional collection of length four consisting of vector bundles. Then the action of the semidirect product $B_{4}\rightthreetimes{\mathbb{Z}}^{4}$ by mutations and shifts on the set of full exceptional collections in ${\mathcal{D}}^{b}(X)$ is transitive and free.
Proof.
Let ${\mathcal{F}}=(F_{0},F_{1},F_{2},F_{3})$ be a full exceptional collection of vector bundles on $X$ and let ${\mathcal{E}}=(E_{0},E_{1},E_{2},E_{3})$ be an arbitrary full exceptional collection in ${\mathcal{D}}^{b}(X)$. Denote by $S_{i}:={F_{i}}|_{Y},S_{i}^{\prime}:={E_{i}}|_{Y}$ the corresponding spherical objects in ${\mathcal{D}}^{b}(Y)$. Proposition 3.13 implies that
$T_{S_{0}}T_{S_{1}}T_{S_{2}}T_{S_{3}}=T_{S_{0}^{\prime}}T_{S_{1}^{\prime}}T_{S_{2}^{\prime}}T_{S_{3}^{\prime}}$
By Theorem 6.5 we are in the context of Lemma 8.5. Hence we conclude that there exists $\sigma\in B_{4}$ such that
$\sigma\cdot\phi({\mathcal{E}})=\phi({\mathcal{F}})=(T_{S_{0}},T_{S_{1}},T_{S_{2}},T_{S_{3}})$
Applying the corresponding sequence of mutations $\sigma\in B_{4}$ to $\mathcal{E}$}, we obtain an exceptional collection $\sigma\cdot{\mathcal{E}}={\mathcal{E}}^{\prime}=(E_{0}^{\prime},\dots,E_{3}^{\prime})$ such that $\phi({\mathcal{E}}^{\prime})=(T_{S_{0}},T_{S_{1}},T_{S_{2}},T_{S_{3}})$. From Lemma 8.4 we deduce that each $E_{i}^{\prime}$, $i=0,\dots,3$ is a shift of a vector bundle. Moreover, $\operatorname{Ext}^{1}(E_{i}^{\prime},E_{i}^{\prime}\otimes\omega_{X}^{-1})=0$ by Remark 8.3, hence applying Theorem 8.1 we obtain the desired transitivity on the set of full exceptional collections.
To prove the freeness it is sufficient to check that the stabiliser of $\mathcal{F}$} is trivial up to shift under the $B_{4}$-action. Observe that the stabiliser up to shift of an exceptional collection $\mathcal{F}$} under the mutation action is contained in the stabiliser of the corresponding tuple $\phi({\mathcal{F}})$ under the Hurwitz action (note that $S\mapsto T_{S}$ kills the shift). On the other hand, by the transitivity proved above, the full $B_{4}$-orbit of $\phi({\mathcal{F}})$ under the Hurwitz action is contained in $\langle T_{S_{0}},T_{S_{1}},T_{S_{2}},T_{S_{3}}\rangle^{4}\subset{\mathcal{G}}^{4}$. Hence by Theorem 7.1(1) this action is in fact the Hurwitz action of $B_{4}$ on $F_{4}^{4}$. It follows from Theorem 3.17 that the stabiliser of $\phi({\mathcal{F}})$ is indeed trivial. ∎
Corollary 8.7.
Every full exceptional collection in ${\mathcal{D}}^{b}(X)$, where $X$ is a threefold with a full exceptional collection of length four consisting of vector bundles, is strong and consists of shifts of vector bundles.
Proof.
Since the property of being strong is preserved under mutations for any smooth projective variety $X$ with $\operatorname{rk}K_{0}(X)=\dim X+1$ ([11], Theorem 2.3), every full exceptional collection is strong by Theorem 8.6. By Proposition 3.4 mutations of exceptional collections of shifted sheaves also consist of shifted sheaves. It is easy to see that in this case collections of shifted locally free sheaves are also preserved by mutations. The conclusion again follows from the transitivity established in Theorem 8.6. ∎
Appendix A Description of the groups $\operatorname{SO}^{+}(q,{\mathbb{Q}})$ and $\operatorname{SO}^{+}(q,{\mathbb{Z}})$
In the Appendix we give an independent proofs for the results of Dolgachev [18] and Kawatani [30] on precise descriptions of the groups $\operatorname{Aut}^{+}{\mathcal{N}}(Y)$ and $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})$ for a K3 surface $Y$ of Picard rank $1$.
Below we put
$\Lambda=\begin{pmatrix}{\mathbb{Z}}&{\mathbb{Z}}\\ \delta{\mathbb{Z}}&{\mathbb{Z}}\end{pmatrix}$.
The group $\Gamma_{0}(\delta)\subset\operatorname{PGL}_{2}^{+}({\mathbb{Q}})$ introduced in §3.4 is the image of $\Lambda^{\ast}\cap\{\det>0\}$ where for a ring $R$ we let $R^{\ast}$ be its group of units.
We now give a local characterization of the Atkin-Lehner elements also introduced in §3.4. For $p$ a prime we put
$\Lambda_{p}=\begin{pmatrix}{\mathbb{Z}}_{p}&{\mathbb{Z}}_{p}\\ \delta{\mathbb{Z}}_{p}&{\mathbb{Z}}_{p}\end{pmatrix}$
where ${\mathbb{Z}}_{p}$ is the localization of $\mathbb{Z}$} at the prime ideal $(p)$.
Lemma A.1.
Let $x\in\operatorname{GL}_{2}^{+}({\mathbb{Q}})$ and let $e$ be an exact divisor of $\delta$. Then $x$ is an Atkin-Lehner element associated to $e$ if and only if
  1. (1)
    For every prime divisor $p$ of $e$ we have
    $x\in\begin{pmatrix}0&-1\\ \delta&0\end{pmatrix}\Lambda^{\ast}_{p}.$
  2. (2)
    For other primes we have
    $x\in\Lambda^{\ast}_{p}.$
Proof.
Let us first prove the $\Leftarrow$ direction. We have
$\begin{pmatrix}0&-1\\ \delta&0\end{pmatrix}\Lambda_{p}=\begin{pmatrix}\delta{\mathbb{Z}}_{p}&{\mathbb{Z}}_{p}\\ \delta{\mathbb{Z}}_{p}&\delta{\mathbb{Z}}_{p}\end{pmatrix}$
and
$\left(\bigcap_{p\mid e}\delta{\mathbb{Z}}_{p}\right)\cap\left(\bigcap_{p\nmid e}{\mathbb{Z}}_{p}\right)=e{\mathbb{Z}}$
so that the hypotheses imply
$x\in\begin{pmatrix}e{\mathbb{Z}}&{\mathbb{Z}}\\ \delta{\mathbb{Z}}&e{\mathbb{Z}}\end{pmatrix}$.
It remains to check that $\det x=e$, or equivalently
$v_{p}(\det x)=\begin{cases}v_{p}(\delta)&\text{if $p\mid e$}\\ 0&\text{otherwise}\end{cases}$
This follows from the hypotheses. The $\Rightarrow$-direction is similar. ∎
The following is the main result in this section.
Proposition A.2 ([18, Remark 7.2.2]).
The isomorphism (4.12) restricts to isomorphisms of groups
(A.1) $\operatorname{PGL}_{2}^{+}({\mathbb{Q}})\cong\operatorname{SO}^{+}(q,{\mathbb{Q}})$
(A.2) $\Gamma^{+}_{0}(\delta)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$
We will use the symbol $\phi$ for a $2\times 2$ matrix representing an element of $\operatorname{PSL}_{2}({\mathbb{R}})$ and the symbol $\Phi$ for the corresponding $3\times 3$-matrix occurring in the formula (4.9). The $\subset$-inclusion in (A.1) is obvious and the $\subset$-inclusion in (A.2) is an immediate verification using the definition of $\Gamma^{+}_{0}(\delta)$.
Hence our task is to prove the $\supset$-inclusions in (A.1) and (A.2).
Assume first $\Phi/\det\phi\in M_{3}({\mathbb{Q}})$. In this case one quickly checks that all defined ratios among the entries of $\phi$ are in $\mathbb{Q}$}. Hence by replacing $\phi$ with a scalar multiple we may and we will assume that $\phi\in M_{2}({\mathbb{Q}})$. From this we obtain (A.1).
It now remains to prove the $\supset$-inclusion in (A.2). We do this locally for each prime.
Lemma A.3.
Assume $\phi\in M_{2}({\mathbb{Q}})$ and $\Phi/\det\phi\in M_{3}({\mathbb{Z}}_{p})$ for some prime number $p$. If $p\nmid\delta$ then, up to a power of $p$, one has $\phi\in\Lambda_{p}^{\ast}$. If $p\mid\delta$ then, up to a power of $p$, one of the following alternatives holds
(A.3) $\phi\in\Lambda_{p}^{\ast}\text{ or }\phi\in\begin{pmatrix}0&-1\\ \delta&0\end{pmatrix}\Lambda_{p}^{\ast}$
Proof.
By multiplying $\phi$ with a general element of $\Lambda^{\ast}_{p}$ we may without loss of generality assume that all entries in $\phi$ are non-zero. By further multiplying $\phi$ with a power of $p$ we may assume that $v_{p}(\det\phi)\in\{0,1\}$. We consider the two cases separately.
  1. (1)
    $v_{p}(\det\phi)=0$. In other words $\det\phi$ is a unit. So we need to check when
    $\begin{pmatrix}d^{2}&2cd&c^{2}/\delta\\ bd&bc+ad&ac/\delta\\ b^{2}\delta&2ab\delta&a^{2}\end{pmatrix}\in M_{3}({\mathbb{Z}}_{p})$
    Looking at $\Phi_{11}$ and $\Phi_{33}$ this immediately gives us $a,d\in{\mathbb{Z}}_{p}$. Put $k=v_{p}(\delta)$. If $k=0$ then looking at $\Phi_{13}$, $\Phi_{31}$ yields $\phi\in M_{2}({\mathbb{Z}}_{p})=\Lambda_{p}$. Since $\det\phi$ is a unit we obtain $\phi\in\Lambda_{p}^{\ast}$. Hence assume $k>0$. By looking at $\Phi_{13}$ one obtains the condition $2v_{p}(c)\geq k$.
    Since adbcad-bc is a unit and $a,d\in{\mathbb{Z}}_{p}$ we must have $v_{p}(b)+v_{p}(c)\geq 0$. Assume first $v_{p}(b)+v_{p}(c)>0$.
    Then $v_{p}(a)=v_{p}(d)=0$. Looking at $\Phi_{21}$ and $\Phi_{23}$ yields $b\in{\mathbb{Z}}_{p}$, $c\in p^{k}{\mathbb{Z}}_{p}$. Hence in this case
    $\phi\in\Lambda_{p}^{\ast}$
    Now assume $v_{p}(b)+v_{p}(c)=0$. From $\Phi_{31}$ we get $2v_{p}(b)+k\geq 0$. Then we obtain that $k$ is even and $v_{p}(c)=-v_{p}(b)=k/2$. We obtain
    $\phi=\begin{pmatrix}yp^{k/2}&vp^{-k/2}\\ wp^{k/2}&xp^{k/2}\end{pmatrix}$
    for $v,w\in{\mathbb{Z}}_{p}^{\ast}$ and $x,y\in{\mathbb{Z}}_{p}$ (the constraint on $\phi_{11}$ comes from $\Phi_{23}$). Hence we have shown
    $\phi\in p^{-k/2}\begin{pmatrix}0&-1\\ \delta&0\end{pmatrix}\Lambda^{\ast}_{p}.$
  2. (2)
    $v_{p}(\det\phi)=1$. So we need to check when
    $\begin{pmatrix}d^{2}/p&2cd/p&c^{2}/(p\delta)\\ bd/p&(bc+ad)/p&ac/(p\delta)\\ b^{2}\delta/p&2ab\delta/p&a^{2}/p\end{pmatrix}\in M_{3}({\mathbb{Z}}_{p}).$
    Looking at $\Phi_{11}$ and $\Phi_{33}$ this immediately gives us $a,d\in p{\mathbb{Z}}_{p}$. Since $v_{p}(ad-bc)=1$}, we get $v_{p}(b)+v_{p}(c)=1$.
    Looking at $\Phi_{13}$ and $\Phi_{31}$ we get
    $2v_{p}(c)\geq k+1$
    $2v_{p}(b)+k-1\geq 0$
    We now get that k=2m1k=2m-1 (in particular, $k\neq 0$ and hence $p\mid\delta$) and $v_{p}(c)=m$, $v_{p}(b)=1-m$. So
    $\phi=\begin{pmatrix}yp^{m}&vp^{-m+1}\\ wp^{m}&xp^{m}\end{pmatrix}$
    with $v,w$ units and $x,y\in{\mathbb{Z}}_{p}$ (as before, the constraints on $\phi_{11}$ and $\phi_{22}$ come from $\Phi_{21}$ and $\Phi_{23}$). Hence
    $\phi\in p^{-m+1}\begin{pmatrix}0&-1\\ \delta&0\end{pmatrix}\Lambda^{\ast}_{p}\qed$
The following corollary finishes the proof of the $\supset$-inclusion in (A.2).
Corollary A.4.
If $\phi\in\operatorname{GL}_{2}^{+}({\mathbb{Q}})$ and $\Phi/\det\phi\in M_{3}({\mathbb{Z}})$ then $\phi$ represents an element of $\Gamma^{+}_{0}(\delta)$.
Proof.
It follows from Lemma A.3 that after multiplying $\phi$ by a scalar we may assume that $\phi\in\Lambda_{p}^{\ast}$ for $p\nmid\delta$ and that (A.3) holds for all primes satisfying $p\mid\delta$. Let $e$ be the exact divisor of $\delta$ whose prime divisors are those for which the second alternative in (A.3) holds. Then it follows from Lemma A.1 that $\phi$ represents an Atkin-Lehner element associated to $e$. ∎
Remark A.5.
In [37, Lemma 1] the author states a similar result as Proposition A.2 for the form $q^{\prime}(r,l,s)=l^{2}-\delta rs$ if $\delta$ is square free. Remarkably, for this modified form the result does not extend to non-square free $\delta$. To see this, observe that in this case we are interested in those $\phi$ for which
$\frac{1}{\det\phi}\begin{pmatrix}d^{2}&2cd&c^{2}\delta\\ bd&bc+ad&ac\delta\\ b^{2}/\delta&2ab/\delta&a^{2}\end{pmatrix}\in{\mathbb{Z}}$
For $\delta=4$ one checks that one may take
$\phi=\begin{pmatrix}a&2b_{1}\\ c_{1}/2&d\end{pmatrix}$
for $a,b_{1},c_{1},d\in{\mathbb{Z}}$, $ad-b_{1}c_{1}=1$. Hence $\operatorname{SO}^{+}(q^{\prime},{\mathbb{Z}})$ is conjugate to $\operatorname{PSL}_{2}({\mathbb{Z}})$ rather than $\Gamma^{+}_{0}(4)$.
Remark A.6.
One checks that (4.11) restricts to an isomorphism
$\operatorname{Aut}^{+}({\mathcal{N}}(Y))/\{\pm 1\}\xrightarrow{\cong}\operatorname{SO}^{+}(q,{\mathbb{Z}})$
so that combining this with (A.2) we obtain an isomorphism
$\operatorname{Aut}^{+}({\mathcal{N}}(Y))/\{\pm 1\}\xrightarrow{\cong}\Gamma^{+}_{0}(\delta)$
Appendix B Understanding $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y)$
To complete our understanding of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ via Theorem 3.20 we need to understand $\operatorname{Aut}^{+}\operatorname{H}^{*}(Y,{\mathbb{Z}})$. We first recall some facts which are specific to the case $d=2$.
Proposition B.1.
[22, Corollary 15.2.12] A K3 surface $Y$ of Picard rank $1$ has a non-trivial automorphism group if and only if it has degree $2$. In the latter case there is a (unique) nontrivial automorphism $\iota\in\operatorname{Aut}(Y)$, which is the covering involution. The involution $\iota$ acts as the identity on ${\mathcal{N}}(Y)$ and as 1-1 on the transcendental lattice $T(Y)={\mathcal{N}}(Y)^{\perp}\subset H^{2}(Y,{\mathbb{Z}})$.
The following result shows that the automorphism $\iota$ is quite remarkable.
Lemma B.2 ([3, Theorem 8.1]).
Let $Y$ be a K3 surface with Picard rank one and degree two. One has:
  1. (1)
    $\iota$ acts trivially on $\operatorname{Stab}^{\dagger}(Y)$;
  2. (2)
    $\iota^{\ast}$ is central in $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$.
  3. (3)
    If $S\in{\mathcal{D}}^{b}(Y)$ is a spherical object then $\iota^{\ast}(S)\cong S$.
For convenience, we give a short self-contained proof.
Proof.
  1. (1)
    Since $\iota$ acts trivially on ${\mathcal{N}}(Y)$, it acts by a deck transformation for $\operatorname{Stab}^{\dagger}(Y)/{\mathcal{P}}^{+}_{0}(Y)$. Hence it must be an element of $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$. This is impossible since $\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$ is torsion free by Theorem 3.21.
  2. (2)
    Let $F\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Then $F\iota^{\ast}F^{-1}$ acts trivially on $\operatorname{Stab}^{\dagger}(Y)$ and so $c:=F\iota^{\ast}F^{-1}\iota^{\ast,-1}$ also acts trivially on $\operatorname{Stab}^{\dagger}(Y)$. On the other hand $c$ acts trivially on $\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$ so $c\in\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)$. By Theorem 3.19 this implies $c=\operatorname{id}$.
  3. (3)
    By Lemma 3.10 and (2) we have $T_{S}=\iota^{\ast}T_{S}\iota^{\ast,-1}=T_{\iota^{\ast}(S)}$ and hence by Lemma 3.11 $\iota^{\ast}S\cong S[l]$ for some $l$. Since $\iota$ is an automorphism, the only possibility is $l=0$. ∎
Proposition B.1 should be contrasted with the following fact.
Lemma B.3.
Let $Y$ be a K3 surface with Picard rank 1 and $d\neq 2$. Then there is no element $\iota$ in $\operatorname{Aut}^{+}H^{\ast}(Y,{\mathbb{Z}})$ which acts as the identity on ${\mathcal{N}}(Y)$ and as 1-1 on $T(Y)$.
Proof.
Assume $\iota$ exists. It follows from the definition that $\iota$ preserves the grading on $H^{\ast}(Y,{\mathbb{Z}})$. In particular it acts on $H^{2}(Y,{\mathbb{Z}})$ such that $\iota(H)=H$ and $\iota|_{T(Y)}=-1$. It follows that $\iota$ is a reflection on $H^{2}(Y,{\mathbb{Z}})$ corresponding to $H$ and so it is given by the standard formula
$(-\iota)(y)=y-\frac{2\langle H,y\rangle}{\langle H,H\rangle}H$
In particular we have $d|2\langle H,y\rangle$ for all $y\in H^{2}(Y,{\mathbb{Z}})$. There can be no $n\neq 1\in{\mathbb{N}}$ which divides $\langle H,y\rangle$ for all $y$ since by unimodularity this would imply $n\mid H$, which is impossible because $H$ is primitive. Hence $d\mid 2$ (and therefore $d=2$). ∎
Recall the following
Proposition B.4 ([22, Corollary 3.3.5]).
Let $Y$ be a K3 surface of odd Picard rank. The only Hodge isometries of the transcendental lattice $T(Y)$ are $\pm{\operatorname{id}}$.
We can now make the following observation.
Lemma B.5.
Let $Y$ be a K3 surface with Picard rank $1$. Let $\operatorname{Aut}^{+}{\mathcal{N}}(Y)$ be the group of isometries of ${\mathcal{N}}(Y)$ preserving the orientation of the positive $2$-planes in ${\mathcal{N}}(Y)_{\mathbb{R}}$. Then the restriction map
(B.1) $\operatorname{Aut}^{+}H^{*}(Y,{\mathbb{Z}})\rightarrow\operatorname{Aut}^{+}{\mathcal{N}}(Y)$
is well-defined. Moreover it is injective for $d>2$ and for $d=2$ its kernel is given by the action of the involution $\iota$ (see Lemma B.1)
Proof.
An element of $\operatorname{Aut}^{+}H^{*}(Y,{\mathbb{Z}})$ acts on the transcendental lattice $T(Y)$ by $\pm\operatorname{id}$, according to Proposition B.4. Since $\pm{\operatorname{id}}$ does not change the orientation of the positive 2-planes in $T(Y)_{\mathbb{R}}$, the map (B.1) is indeed well-defined.
If $\phi$ is a non-trivial element in the kernel of (B.1) then it is the identity on ${\mathcal{N}}(Y)$, and it acts as 1-1 on $T(Y)$. Invoking Proposition B.1 and Lemma B.3 finishes the proof. ∎
Our next aim is to characterize when (B.1). This is tackled by the following result.
Proposition B.6 ([18, Theorem 7.1]).
The image of the map
$\operatorname{Aut}^{+}H^{*}(Y,{\mathbb{Z}})\to\operatorname{Aut}^{+}{\mathcal{N}}(Y)\to\operatorname{SO}^{+}(q,{\mathbb{Z}})\cong\Gamma_{0}^{+}(\delta)$
is equal to $F(\delta)$.
Proof.
Recall that ${\mathcal{N}}(Y)\cong\operatorname{H}^{0}(Y,{\mathbb{Z}})\oplus\operatorname{Pic}(Y)\oplus\operatorname{H}^{4}(Y,{\mathbb{Z}})$. After identifying $\operatorname{H}^{0}(Y,{\mathbb{Z}})\cong{\mathbb{Z}}\cong\operatorname{H}^{4}(Y,{\mathbb{Z}})$, an element $\Phi\in\operatorname{Aut}^{+}({\mathcal{N}}(Y))$ can be identified with an element of
(B.2) $\operatorname{End}({\mathbb{Z}}\oplus\operatorname{Pic}(Y)\oplus{\mathbb{Z}})=\begin{pmatrix}{\mathbb{Z}}&\operatorname{Pic}(Y)^{\vee}&{\mathbb{Z}}\\ \operatorname{Pic}(Y)&\operatorname{End}(\operatorname{Pic}(Y),\operatorname{Pic}(Y))&\operatorname{Pic}(Y)\\ {\mathbb{Z}}&\operatorname{Pic}(Y)^{\vee}&{\mathbb{Z}}\end{pmatrix}$
Our task is to understand when such an element can be lifted to an element of
(B.3) $\operatorname{End}(\operatorname{H}^{0}(Y,{\mathbb{Z}})\oplus\operatorname{H}^{2}(Y,{\mathbb{Z}})\oplus\operatorname{H}^{4}(Y,{\mathbb{Z}}))=\begin{pmatrix}{\mathbb{Z}}&\operatorname{H}^{2}(Y,{\mathbb{Z}})^{\vee}&{\mathbb{Z}}\\ \operatorname{H}^{2}(Y,{\mathbb{Z}})&\operatorname{End}(\operatorname{H}^{2}(Y,{\mathbb{Z}}),\operatorname{H}^{2}(Y,{\mathbb{Z}}))&\operatorname{H}^{2}(Y,{\mathbb{Z}})\\ {\mathbb{Z}}&\operatorname{H}^{2}(Y,{\mathbb{Z}})^{\vee}&{\mathbb{Z}}\end{pmatrix}$
which acts as $\pm 1$ on $T(Y)$. Note that by Lemma B.5 such an extension is unique for $d>2$ and for $d=2$ there is a ${\mathbb{Z}}/2{\mathbb{Z}}$ ambiguity which is inconsequential for what follows.
The only troublesome part of (B.3) is the middle column, since filling its entries involves extending a map with source $\operatorname{Pic}(Y)$ to a map with source $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ and $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ is not the direct sum of $\operatorname{Pic}(Y)$ and $T(Y)$ (it cannot be since $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ is unimodular). Now recall
  1. (1)
    The kernel of $\operatorname{Aut}^{+}{\mathcal{N}}(Y)\rightarrow\operatorname{SO}^{+}(q,{\mathbb{Z}})$ is $\{\pm 1\}$ by Remark A.6.
  2. (2)
    The identification $\operatorname{SO}^{+}(q,{\mathbb{Z}})\cong\Gamma^{+}_{0}(\delta)$ is given by an explicit formula (4.9).
  3. (3)
    By definition $\Gamma^{+}_{0}(\delta)$ is the collection of (images) of Atkin-Lehner elements in $\operatorname{PSL}_{2}({\mathbb{R}})$ (see Definition 3.30).
Hence our mission is to determine when an Atkin-Lehner element can be lifted to $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$. Let
$\begin{pmatrix}ea&b\\ \delta c&ed\end{pmatrix}\in\Gamma_{0}^{+}(\delta)$
be an arbitrary Atkin-Lehner element. Then up to sign, and after identifying $\operatorname{Pic}(Y)={\mathbb{Z}}H$, the corresponding middle column of (B.2) is
$\begin{pmatrix}(2\delta cd)H^{\vee}\\ \left({\displaystyle\frac{\delta}{e}}bc+ead)(H^{\vee}\otimes H\right)\\ (2\delta ab)H^{\vee}\end{pmatrix}$
where $H^{\vee}$ denotes the dual basis element to $H$. Since $\langle H,H\rangle=2\delta$, the map $(2\delta cd)H^{\vee}:{\mathbb{Z}}H\rightarrow{\mathbb{Z}}$ extends to $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ to a map which is zero on $T(Y)$ via $cd\langle H,$-$\rangle$. Similarly, $(2\delta ab)H^{\vee}$ extends to $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ via $ab\langle H,$-$\rangle$.
It remains to consider the middle entries of (B.2) and (B.3). We analyze when an element of $\beta(H^{\vee}\otimes H)\in{\mathbb{Z}}$ extends to map $f:\operatorname{H}^{2}(Y,{\mathbb{Z}})\rightarrow\operatorname{H}^{2}(Y,{\mathbb{Z}})$ which is $\pm\operatorname{id}$ on $T(Y)$. We assume first that it extends as $\operatorname{id}$. Then $f$ has the form $x\mapsto x+\alpha\langle H,x\rangle H$ where $\alpha\in{\mathbb{Z}}$ because of the unimodularity of $\operatorname{H}^{2}(Y,{\mathbb{Z}})$.
Expressing $f(H)=\beta H$ yields $1+2\delta\alpha=\beta$, so $(2\delta)|(\beta-1)$. Applying this with $\beta=\frac{\delta}{e}bc+ead$ and recalling that $ead-\frac{\delta}{e}bc=1$ by the definition of an Atkin-Lehner element, we see that this implies $e|bc$, but this contradicts $ead-\frac{\delta}{e}bc=1$ unless $e=1$. Similarly, $f$ extends to $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ as 1-1 on $T(Y)$ if and only if $e=\delta$. ∎
Corollary B.7.
$\operatorname{Aut}^{+}(\operatorname{H}^{\ast}(Y,{\mathbb{Z}}))\rightarrow\operatorname{Aut}^{+}({\mathcal{N}}(Y))$ is surjective if and only if $\delta$ is a prime power.
Proof.
We have surjectivity if and only if $\operatorname{Aut}^{+}(\operatorname{H}^{\ast}(Y,{\mathbb{Z}}))/\{\pm 1\}\rightarrow\operatorname{Aut}^{+}({\mathcal{N}}(Y))/\{\pm 1\}$ is surjective. I.e. if an only if $F(\delta)=\Gamma^{+}_{0}(\delta)$ which is equivalent to $\delta$ being a prime power. ∎
We now obtain a version of (3.6) in which everything is explicit
Theorem B.8.
Let $N$ be the subgroup of $\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ generated by
$\begin{cases}[1]&\text{if $d\neq 2$}\\ [1],\iota^{\ast}&\text{if $d\neq 2$}\end{cases}$
(recall that by Lemma B.2(2) $\iota^{\ast}$ is central) and let $\bar{N}$ be its image in $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})$. I.e.
$\bar{N}=\begin{cases}\{\pm 1\}&\text{if $d\neq 2$}\\ \{\pm 1,\pm\iota\}&\text{if $d\neq 2$}\end{cases}$
There is a short exact sequence
(B.4) $1\rightarrow\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]\rightarrow\operatorname{Aut}{\mathcal{D}}^{b}(Y)/N\rightarrow F(\delta)\rightarrow 1$
Proof.
From Theorem 3.20 we obtain in fact an exact sequence
$1\rightarrow\operatorname{Aut}^{0}{\mathcal{D}}^{b}(Y)/[2]\rightarrow\operatorname{Aut}{\mathcal{D}}^{b}(Y)/N\rightarrow\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})/\bar{N}\rightarrow 1$
Now we observe that there is an injection $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y,{\mathbb{Z}})/\bar{N}\hookrightarrow\operatorname{Aut}^{+}{\mathcal{N}}(Y)/\{\pm 1\}$ (Lemma B.5) and an isomorphism $\operatorname{Aut}^{+}{\mathcal{N}}(Y)/\{\pm 1\}\cong\Gamma^{+}_{0}(\delta)$ (Remark A.6). The image under the composition of $\operatorname{Aut}^{+}H^{\ast}(Y,{\mathbb{Z}})$ was identified with $F(\delta)$ in Proposition B.6. ∎